### v2

# Stochastic Hessian Fittings on Lie Groups

Abstract
This paper studies the fitting of Hessian or its inverse for stochastic optimizations using a Hessian fitting criterion from the preconditioned stochastic gradient descent (PSGD) method, which is intimately related to many commonly used second order and adaptive gradient optimizers, e.g., BFGS, Gaussian-Newton and natural gradient descent, AdaGrad, etc. Our analyses reveal the efficiency and reliability differences among a wide range of preconditioner fitting methods, from closed-form to iterative solutions, using Hessian-vector products or stochastic gradients only, with Hessian fittings in the Euclidean space, the manifold of symmetric positive definite (SPL) matrices, or a variety of Lie groups. The most intriguing discovery is that the Hessian fitting itself as an optimization problem is strongly convex under mild conditions on a specific yet general enough Lie group. This discovery turns Hessian fitting into a well behaved optimization problem, and facilitates the designs of highly efficient and elegant Lie group sparse preconditioner fitting methods for large scale stochastic optimizations.
Index Terms:
Stochastic gradient descent (SGD), Hessian-vector product, Hessian fitting, Lie group, strong convexity.
I Introduction
This paper considers the problem of Hessian fitting for second order stochastic optimizations where the objective functions are typically defined as an expectation, say $f(\theta)=E_{z}[\ell(\theta,z)]$ with $\theta\in\mathbb{R}^{m}$ being the parameters to be optimized, and $z$ a random vector that can be sampled to evaluate an unbiased estimate of $f(\theta)$. We generally assume that $f(\theta)$ is second-order differentiable such that a preconditioner $P\succ 0$, say the inverse of the Hessian of $f(\theta)$ when it is convex, can be used to accelerate the convergence of the following iteration,
$\theta_{t+1}=\theta_{t}-\mu_{t}P_{t}\,{\partial f_{t}(\theta)}/{\partial\theta}\left.\right|_{\theta=\theta_{t}}$ (1)
to an optimal solution for $\theta$, where subscript $t$ is the time index, and $f_{t}(\theta)$ a sampled noisy estimation of $f(\theta)$ at step $t$. To simplify the notations, let us ignore the time index $t$ when doing so causing no misunderstanding, and use $g={\partial f(\theta)}/{\partial\theta}$ and $h={\partial(v^{T}g)}/{\partial\theta}=Hv$ to denote the stochastic gradient and Hessian-vector product associated with a vector $v\sim\mathcal{N}(0,I)$ at a certain iteration step, respectively, where $H={\partial^{2}f(\theta)}/{(\partial\theta^{T}\partial\theta)}$ is the Hessian and we typically do not assume its definiteness. Then, with the following preconditioner estimation criterion proposed in [17],
$c(P;v,h)=h^{T}Ph+v^{T}P^{-1}v$ (2)
we can derive most of the commonly used preconditioners as listed in Table 1 [18].
  1. 1.
    When the curvature condition $v^{T}h>0$ is met, solution $Phh^{T}P=vv^{T}$ reduces to the secant equation $Ph=v$ , which is used for deriving the classic BFGS and limited-memory BFGS (L-BFGS) like quasi-Newton methods via iterative Hessian fittings [1]. These off-the-shelf methods are not popular for stochastic optimizations as they do not take gradient noises into considerations.
  2. 2.
    Closed-form solution $P=(H^{2})^{-0.5}$ reduces to the Newton’s method when $H\succ 0$, i.e., $P=H^{-1}$. Diagonal form of this solution is particularly popular in machine learning due to its simplicity, e.g., ESGD [12], AdaHessian [13], and Sophia [14] to name a few.
  3. 3.
    When $g={\partial\ell(\theta,z)}/{\partial\theta}$ is a sample-wise gradient for each $z$, closed-form solution $P^{-2}=E_{z}[gg^{T}]$ reduces to the (empirical) Fisher information matrix, whose inverse is used in the Gaussian-Newton and natural gradient methods [3] as preconditioners. Its Kronecker factorized form is a popular approximation for the learning of matrix parameters [5, 6]. The Fisher information matrix and Hessian are closely related when $\ell(\cdot)$ is a negative logarithm likelihood (NLL) loss from the exponential family [2].
  4. 4.
    Closed-form solution $P_{t}=\left(\sum_{i=1}^{t}g_{i}g_{i}^{T}\right)^{-0.5}$ reduces to the preconditioner used in the AdaGrad [4], a large family of optimizers including many popular examples, e.g., the RMSProp [10] and Adam(W) [11] that use diagonal preconditioners, and a few methods with name Shampoo the most well known [6, 7, 8, 9] that use Kronecker factorized approximate closed-form solutions.
Except for the trivial case of diagonal preconditioners, all the last three closed-form solutions for $P$ in Table I require matrix or matrix square root inverse operations, which could cause numerical difficulties in practice [8]. Furthermore, exact closed-form solutions are seldom available for sparse Hessian estimations. The iterative Hessian fitting method of BFGS requires that $v^{T}h$ is well above zero, which can be hardly met in stochastic optimizations. The preconditioned stochastic gradient descent (PSGD) fits $P$ iteratively on a variety of Lie groups [17, 18, 19], and seems successfully avoid the numerical difficulties in closed-form solutions and also lift the limitations of BFGS. Recent empirical results have further confirmed PSGD’s excellent performance on a handful of large scale machine learning related stochastic optimizations [20, 21]. Still, there lacks a deeper theoretical understandings of the pros and cons of the preconditioners listed in Table I, disregarding their certain practical difficulties like computational complexities and numerical difficulties with finite precision arithmetic. One major contribution of this paper is to show that the problem of Hessian fitting by minimizing criterion (2) converges linearly to the optimal solution on the manifold of symmetric positive definite (SPD) matrices, and furthermore, is strongly convex on a certain Lie group, see Proposition 5 in Section 3.1, while those closed-form solutions only converge sublinearly. Another significant contribution is the invention of a L-BFGS like low-rank approximation (LRA) version of PSGD  [^1] This work has been presented in the NeurIPS 2022 HOOML workshop, a non-archived channel. . Lastly, the insights brought up by our studies also help to improve previous PSGD implementations [17, 18, 19] in several aspects.
Notations in the paper are fairly standard. For example, $A\succ 0$ means that $A$ is a SPD matrix, $E_{v}[\cdot]$ takes expectation with respect to a certain random variable $v$, $\mathcal{N}(0,I)$ denotes the standard normal distribution, $A^{T}$ is the transpose of $A$, $A^{0.5}$ denotes the principal square root of $A$, $A^{-2T}$ is the shorthand for $(A^{-2})^{T}$, $\|A\|_{2}$ is the spectral norm of $A$, ${\rm tr}(A)$ takes the trace of $A$, $\lambda_{\min}(A)$ takes the minimum eigenvalue of $A$, ${\rm vec}(A)$ vectorizes matrix $A$ by stacking its columns successively, GL$(n,\mathbb{R})$ stands for the general linear group of $n\times n$ invertible matrices, and $\otimes$ denotes Kronecker product. We say that $A$ and $B$ commutes if $[A,B]=AB-BA=0$. For a symmetric matrix $A\in\mathbb{R}^{n\times n}$, we only need to vectorize either its upper or lower triangular part, and denote this operation as ${\rm vech}(A)$. Clearly, ${\rm vec}(A)=S\,{\rm vech}(A)$, where $S\in\mathbb{R}^{n^{2}\times[0.5n(n+1)]}$ is a sparse full column rank matrix with elements being 00 or $1$. By convection, we use capital letters for matrices, and lower case letters for vectors and scalars.
TABLE I: Variations of criterion (2) for preconditioner fitting, where $v$ is a dummy variable in 2), 3) and 4) as $E_{v}[v^{T}P^{-1}v]={\rm tr}(P^{-1})$.
Criterion Solution for $P$
1) $c(P;v,h)$ $Phh^{T}P=vv^{T}$
2) $E_{v}[c(P;v,h)]$ $P^{-2}=H^{2}$
3) $E_{v,z}[c(P;v,g)]$ $P^{-2}=E_{z}[gg^{T}]$
4) $\sum_{i=1}^{t}E_{v_{i}}\left[c(P;v_{i},g_{i})\right]$ $P_{t}^{-2}=\sum_{i=1}^{t}g_{i}g_{i}^{T}$
II Candidates for Stochastic Hessian Fitting
II-A The Problem Setups
We are to fit the Hessian or its inverse sequentially, i.e., updating the Hessian estimate once given every pair of vector and its associated Hessian-vector product, $(v,h)$, with model,
$h=Hv+\epsilon$ (3)
where $H\in\mathbb{R}^{n\times n}$ is the unknown Hessian matrix to be estimated, and $\epsilon$ denotes any possible modeling errors, e.g., truncation and quantization errors when the Hessian-vector product $h$ is approximated with a finite-difference method, or the randomness error when a stochastic yet unbiased proxy of the target function $f(\theta)=E_{z}[\ell(\theta,z)]$ is adopted for Hessian-vector product evaluation, or the structural errors caused by using a certain form of sparse matrices to fit dense Hessians. We assume that $\epsilon$ is a random variable independent of $v$.
We choose to update $P$ by minimizing the preconditioner fitting criterion (2), whose expectation with respect to $v$ is,
$c(P)=$ $E_{v\sim\mathcal{N}(0,I)}[c(P;v,h)]$
$=$ ${\rm tr}\left(PE[hh^{T}]+P^{-1}E[vv^{T}]\right)$
$=$ ${\rm tr}\left(P\left(H^{2}+E[\epsilon\epsilon^{T}]\right)+P^{-1}\right)$ (4)
Then, the optimal SPD solution minimizing $c(P)$ can be shown to be [17]
$P=\left(H^{2}+E[\epsilon\epsilon^{T}]\right)^{-0.5}$ (5)
by solving equation ${\partial c(P)}/{\partial P}=0$ for $P$. One also can replace the $P$ in (2) with $P^{-1}$ to obtain another criterion for fitting the Hessian directly. Here, we only consider the fitting criterion (2) since the intention is to use $P$ as a preconditioner for accelerating the SGD iteration in (1).
As revealed by (5), the error term $\epsilon$ in (3) helps to damp the preconditioner estimation and has been shown to regularize the optimizer in a perfectly balanced way [17]. Use of $(H^{2})^{-0.5}$ as a preconditioner directly could severely amplify the gradient noises [17]. Definiteness of the Hessian $H$ is important for the optimization of $f(\theta)$, but not interested to our setups since both (4) and (5) suggest that the signs of $\lambda(H)$ do not matter to the finial solutions for $P$. Hence, without loss of generality, we always assume that $H$ is an invariant positive definite matrix for our studies. Just bear in mind that except for the BFGS, other preconditioners derived from (2) are always valid regardless of the definiteness of $H$.
We will further ignore the term $\epsilon$ most of the time to simplify our performance analyses. Still, note that $c(P)$ always has form $c(P)={\rm tr}(PA+P^{-1})$, where $A=H^{2}+E[\epsilon\epsilon^{T}]$ with pair $(v,h)$, and $A=E_{z}[gg^{T}]$ or $\sum_{i}[g_{i}g_{i}^{T}]$ with pair $(v,g)$. Thus, properties proved for one case can be easily transferred to the other cases.
Criterion (2) is a well defined function for any $P$ with $\det(P)\neq 0$, i.e., the group GL$(n,\mathbb{R})$. For our purpose, we further require $P\succ 0$ such that it is a valid preconditioner for the SGD iteration in (1). This symmetry constraint, i.e., $P=P^{T}$, will slightly complicate the calculations of the derivatives of $c(P;v,h)$ with respect to $P$. We will bring up this subtleness explicitly when it matters. From now on, by default, the gradient and Hessian always refer to the ones of $c(P;v,h)$ or $c(P)=E_{v}[c(P;v,h)]$ with respect to $P$. Gradient descent (GD) or SGD also refers to the iterations for updating $P$, not $\theta$, if not further clarified.
II-B Second Order Approximation of Criterion (2)
Let $dP$ denote a small perturbation around $P$. We are to expand the fitting loss in (2) up to its second order terms of $dP$. To begin with, we expand $(P+dP)^{-1}$ around $P^{-1}$ as,
$\quad dP^{-1}=(P+dP)^{-1}-P^{-1}$
$=(I-P^{-1}dP+P^{-1}dP\,P^{-1}dP+\mathcal{O}((dP)^{3}))P^{-1}-P^{-1}$
$=-P^{-1}dP\,P^{-1}+P^{-1}dP\,P^{-1}dP\,P^{-1}+\mathcal{O}((dP)^{3})$
where $\mathcal{O}((dP)^{3})$ includes the third and higher order terms of $dP$ that are negligible when $\|dP\|$ is small enough. Now, with the above expression, we can expand (2) as below,
$dc(P;v,h)=$ $h^{T}dP\,h+v^{T}dP^{-1}\,v$
$=$ $h^{T}dP\,h-v^{T}P^{-1}dP\,P^{-1}v$
$+v^{T}P^{-1}dP\,P^{-1}dP\,P^{-1}v$ (6)
where we have suppressed the term $\mathcal{O}((dP)^{3})$ to simplify our equation. Clearly, the $dc(P;v,h)$ in (6) becomes a quadratic function of $dP$ after truncating higher order terms. To maximize the reduction of $c(P;v,h)$, we can solve for the optimal $dP$ by letting $dc(P;v,h)/dP=0$, i.e., the set of equations,
$hh^{T}-P^{-T}vv^{T}P^{-T}+P^{-T}vv^{T}P^{-T}dP^{T}\,P^{-T}$
$+P^{-T}dP^{T}\,P^{-T}vv^{T}P^{-T}=0$ (7)
where we do not assume $P=P^{T}$ yet to simplify the calculation of derivatives. From (7), it is ready to identify the gradient of $c(P;v,h)$ with respect to $P$ as below,
$\frac{\partial c(P;v,h)}{\partial P}=hh^{T}-{P}^{-T}vv^{T}{P}^{-T}$ (8)
Note that (8) gives the steepest ascent direction of $c(P;v,h)$ in the space $\mathbb{R}^{n\times n}$. With constraint $P=P^{T}$, we should double the off-diagonal elements of the gradient in (8) to obtain the steepest ascent direction in the space $\mathbb{R}^{0.5n(n+1)}$. We can choose either form for the update of $P$ with GD or SGD. But, we do need to calculate the Hessian with respect to $P$ in space $\mathbb{R}^{0.5n(n+1)}$. We typically take the expectation of criterion (2), i.e., the $c(P)=E_{v\sim\mathcal{N}(0,I)}[c(P;v,h)]$ given in (4), for the performance analyses. We have the following convex property of $c(P)$.
Proposition 1: Criterion $c(P)$ is a convex function of $P$ on set $\{P|P\succ 0\}$.
Proof: From (6), we can rewrite the second order term of $dc(P)$ as
$\quad E_{v\sim\mathcal{N}(0,I)}\left[v^{T}P^{-1}dP\,P^{-1}dP\,P^{-1}v\right]$
$={\rm tr}\left(E_{v}\left[P^{-1}dP\,P^{-1}dP\,P^{-1}vv^{T}\right]\right)$
$={\rm tr}\left(P^{-1}dP\,P^{-1}dP\,P^{-1}\right)$
$=\left({\rm vec}\left(\left(P^{-1}dP\right)^{T}\right)\right)^{T}{\rm vec}\left(P^{-1}dP\,P^{-1}\right)$
$=\left({\rm vec}\left(dP\,P^{-1}\right)\right)^{T}{\rm vec}\left(P^{-1}dP\,P^{-1}\right)$
$=\left(\left(P^{-1}\otimes I\right){\rm vec}(dP)\right)^{T}\left(\left(P^{-1}\otimes P^{-1}\right){\rm vec}(dP)\right)$
$=\left({\rm vec}(dP)\right)^{T}\left(P^{-1}\otimes I\right)\left(P^{-1}\otimes P^{-1}\right){\rm vec}(dP)$
$=\left({\rm vec}(dP)\right)^{T}\left(P^{-2}\otimes P^{-1}\right){\rm vec}(dP)$
$=\left({\rm vech}(dP)\right)^{T}S^{T}\left(P^{-2}\otimes P^{-1}\right)S\,{\rm vech}(dP)$ (9)
where we have used the relation ${\rm tr}(A^{T}B)=\left({\rm vec}(A)\right)^{T}{\rm vec}(B)$ to arrive the fourth line from the third one, ${\rm vec}(ABC)=(C^{T}\otimes A){\rm vec}(B)$ from the fifth to the sixth line, $(AB)\otimes(CD)=(A\otimes C)(B\otimes D)$ from the third to the last to the next line, and definition ${\rm vec}(A)=S\,{\rm vech}(A)$ for $A\succ 0$ to get the last line. Since $P^{-2}\otimes P^{-1}\succ 0$ and $S$ has full column rank, the Hessian with respect to $P$, i.e.,
$S^{T}\left(P^{-2}\otimes P^{-1}\right)S$
due to the last line of (9), is positive definite. $\square$
Note that following the proof of Proposition 1, we also can show that $c(P;v,h)$ is convex as well. Actually, this convexity property is independent of the definiteness of $H$. In general, the quadratic term in (6) can be negative without constraint $dP=dP^{T}$ even when $P\succ 0$. Thus, the symmetry constraint is essential for the convexity of $c(P)$.
II-C Hessian Fitting in the Euclidean Space
II-C1 Gradient Descent
One obvious choice is to update $P$ with the gradient given in (8). Let us start from an initial guess $P_{0}\succ 0$, and update $P$ at step $t$ with SGD as
$P_{t+1}=P_{t}-\mu(h_{t}h_{t}^{T}-{P}_{t}^{-1}v_{t}v_{t}^{T}{P}_{t}^{-1})$ (10)
where $\mu$ is a small enough positive step size, $(v_{t},h_{t})=(v_{t},Hv_{t})$ with $v_{t}\sim\mathcal{N}(0,I)$ is a pair of vector and its associated Hessian-vector product at step $t$. Note that $P_{t}$ is always symmetric as long as $P_{0}$ is.
From the proof of Proposition 1, we see that the Hessian of $c(P)$ with respect to $P$ is neither lower nor upper tightly bounded. Thus, it is generally not easy to characterize the convergence rate of iteration (10). Still, following the regret analysis [4, 11], we can show that with a proper constant step size $\mu$, the sequence given by (10) converges to $H^{-1}$ at least sublinearly when $H\succ 0$. We only sketch the key steps here. Assuming that $\mu$ is small enough such that we always have $P_{t}\succ 0$ as long as $P_{0}\succ 0$. Then, via the convexity of $c(P;v,h)$, we can show that
$\quad c(P_{t};v_{t},h_{t})-c(H^{-1};v_{t},h_{t})$
$\leq\frac{\left\|P_{t}-H^{-1}\right\|_{F}^{2}-\left\|P_{t+1}-H^{-1}\right\|_{F}^{2}}{2\mu}+\frac{\mu}{2}\left\|\frac{\partial c(P_{t};v_{t},h_{t})}{\partial P_{t}}\right\|_{F}^{2}$
Then, we take expectation on both sides of the above equation with respect to $v$, and sum over the time index $t$ to have
$c\left(\sum_{i=1}^{t}P_{i}/t\right)-c(H^{-1})=\mathcal{O}(1/\sqrt{t})$
by setting $\mu\propto 1/\left(\sqrt{t}\,\max_{1\leq i\leq t}E_{v_{i}}\left[\left\|\frac{\partial c(P_{i};v_{i},h_{i})}{\partial P}\right\|_{F}\right]\right)$. Still, the regret analysis does not exclude possible better convergence results of (10) with time-varying step sizes.
II-C2 The Newton’s Method
Another choice is to take the expectation on the left side of (7) to have
$H^{2}-P^{-2}+P^{-2}dP\,P^{-1}+P^{-1}dP\,P^{-2}=0$ (11)
and then solve for the optimal $dP$ either directly as
$\>\>{\rm vec}(dP)=\left(P^{-1}\otimes P^{-2}+P^{-2}\otimes P^{-1}\right)^{-1}{\rm vec}\left(P^{-2}-H^{2}\right)$
$=\left(I\otimes P^{-1}+P^{-1}\otimes I\right)^{-1}(P\otimes P){\rm vec}\left(P^{-2}-H^{2}\right)$
$=\left(I\otimes P^{-1}+P^{-1}\otimes I\right)^{-1}{\rm vec}\left(I-PH^{2}P\right)$ (12)
or via the matrix sign function iteratively [22], which is clear after rewriting (11) into the following familiar continuous Lyapunov function form
$P^{-1}dP+dP\,P^{-1}=I-PH^{2}P$
where $H^{2}$ is given or estimated from $H^{2}=E[hh^{T}]$ by replacing $E[\cdot]$ with sample averages. Eq. (12) or the method in [22] gives a numerically stable but costly solution for $dP$.
But, if we assume that $P$ and $dP$ commutes such that $P^{-2}dP\,P^{-1}=P^{-1}dP\,P^{-2}$, then (11) immediately gives the optimal $dP$ as,
$dP=-0.5P(H^{2}-P^{-2})P^{2}=0.5(P-PH^{2}P^{2})$ (13)
Now, (13) suggests the following Newton’s method for updating $P$,
$P_{t+1}=1.5P_{t}-0.5P_{t}H^{2}P_{t}^{2}$ (14)
Proposition 2: Starting from an initial guess $P_{0}\succ 0$ that commutes with $H$ and satisfies condition $20<\lambda(HP)<(\sqrt{17}-1)/$ 2, the Newton’s method in (14) converges to $H^{-1}$ quadratically.
Proof: First, let us show that $P_{t+1}$ commutes with $H$ if $P_{t}$ does. With (14), we have
$P_{t+1}H-HP_{t+1}$
$=$ $1.5(P_{t}H-HP_{t})-0.5(P_{t}H^{2}P_{t}^{2}H-HP_{t}H^{2}P_{t}^{2})$
$=$ $-0.5(P_{t}H^{3}P_{t}^{2}-P_{t}H^{3}P_{t}^{2})=0$
Hence, $P_{t}$ always commutes with $H$ as long as $P_{0}$ does. Also, we can easily show that $P_{t}$ is symmetric as long as $P_{0}$ is.
Now, let us define $R_{t}=HP_{t}-I$. Similarly, we can show that $P_{t}$, $H$ and $R_{t}$ all commute with each other and are symmetric. Then, by (14), we have
$R_{t+1}=$ $1.5HP_{t}-0.5HP_{t}H^{2}P_{t}^{2}-I$
$=$ $1.5HP_{t}-0.5(HP_{t})^{3}-I$
$=$ $1.5(R_{t}+I)-0.5(R_{t}+I)^{3}-I$
$=$ $-0.5(R_{t}+3I)R_{t}^{2}$ (15)
We require
$\|0.5(R_{t}+3I)R_{t}\|_{2}<1$
to make (15) a contraction mapping such that $\|R_{t+1}\|_{2}<\|R_{t}\|_{2}$, which suggests that the eigenvalues of $P_{t}H$ should be bounded by
$0<\lambda(P_{t}H)<\frac{\sqrt{17}-1}{2}$ (16)
Then, with the bounds given in (16) and the last line of (15), we can show the following quadratic convergence,
$\frac{\|R_{t+1}\|_{2}}{\|R_{t}\|_{2}^{2}}=0.5\|R_{t}+3I\|_{2}<\frac{\sqrt{17}+3}{4}$
Note that the lower bound of $\lambda(P_{t}H)$ in (16) already implies that $H$ cannot be singular such that $H^{-1}$ exists. $\square$
It is trivial to choose a starting point $P_{0}$ that commutes with $H$, e.g., a $P_{0}\propto I$. Indeed, starting from a $P_{0}\propto I$, we see that $P_{t}$ is a linear combination of terms in $\{I,H,\ldots,H^{2t+2}\}$ and thus commutes with $H$. However, in practice, the Newton’s method given in (14) is vulnerable to round-off errors that eventually break the commuting property between $P_{t}$ and $H$.
Another related method is the Newton-Schulz iteration [23], which also can be used to iteratively calculate $(H^{2})^{-0.5}$. We seldom use these Newton’s methods in practice. We present them here mainly for theoretical interests, especially their connections to the matrix square root calculations [22, 23].
II-C3 A Few Closed-Form Solutions
For the sake of completeness, let us also briefly discuss the convergence rate of a few commonly used closed-form solutions. The most straightforward choice is $P_{t}=\left(\left(P_{0}^{-2}+\sum_{i=1}^{t}h_{i}h_{i}^{T}\right)/(t+1)\right)^{-0.5}$ used in methods from the AdaGrad family [4], which can be put in a more fancier form as below
$P_{t+1}=\left(\frac{t+1}{t+2}P_{t}^{-2}+\frac{1}{t+2}h_{t}h_{t}^{T}\right)^{-0.5}$ (17)
As $h_{t}=Hv_{t}$ and $v_{t}\sim\mathcal{N}(0,I)$, we have
$P_{t}=$ $\left(H\left(P_{0}^{-2}/(t+1)+\sum_{i=1}^{t}v_{i}v_{i}^{T}/(t+1)\right)H\right)^{-0.5}$
$=$ $\left(H\left(I+\mathcal{O}(1/t)\right)H\right)^{-0.5}$
$=$ $H^{-1}+\mathcal{O}(1/t)$
Hence, the closed-form solution given by (17) converges to $H^{-1}$ sublinearly with rate $\mathcal{O}(1/t)$ when $H\succ 0$.
Another closed-formed solution can be obtained from (7) by solving for $P$ directly after setting $dP=0$. Specifically, (7) suggests that a stationary point $P$ satisfies
$P\left(\sum_{i=1}^{t}h_{i}h_{i}^{T}\right)P-\sum_{i=1}^{t}v_{i}v_{i}^{T}=0$ (18)
Then, we can directly solve the above Riccati equation for $P$ when both $\sum_{i=1}^{t}v_{i}v_{i}^{T}$ and $\sum_{i=1}^{t}h_{i}h_{i}^{T}$ are not singular, see the proof of Proposition 3 in [17]. Clearly, (18) reduces to (17) for $t\rightarrow\infty$. But, this solution could be exact even with a finite $t$. Similar to the Newton’s method, this closed-form solution is mainly for theoretical interests.
One more family of solutions worth to mention here are the BFGS like methods that modify $P_{t}$ with certain rank-2 updates such that $P_{t+1}h_{t}=v_{t}$ [1]. They are essentially iterative methods for Hessian fittings, but with a few heuristic closed-form update rules for $P$. These methods have two major limitations: the pair $(v_{t},h_{t})$ needs to satisfy the curvature condition $v_{t}^{T}h_{t}>0$; they diverge easily with noisy Hessian-vector products as inputs. Note that generally, these heuristic methods could diverge, or might not converge to $\left(H^{2}\right)^{-0.5}$.
II-D Hessian Fitting on the Manifold of SPD
One possible way to define a local coordinate for $P$ on the manifold of SPD is to let $dP$ be
$dP=P\mathcal{E}+\mathcal{E}P^{T}$ (19)
where $\mathcal{E}$ is a sufficiently small matrix, but not necessarily symmetric in general. Clearly, $dP$ is symmetric if $\mathcal{E}$ is, regardless of the symmetry property of $P$. On the other hand, we can show that $\mathcal{E}$ is symmetric if $P\succ 0$.
Proposition 3: Eq. (19) induces a new coordinate system for the manifold $\{P|P\succ 0\}$.
Proof: We are to show that (19) defines a bijective mapping between $dP$ and $\mathcal{E}$ when $P\succ 0$, which is partially done after we vectorize both sides of (19) to connect $dP$ to $\mathcal{E}$ as,
${\rm vec}(\mathcal{E})=(I\otimes P+P\otimes I)^{-1}{\rm vec}(dP)$ (20)
Still, we need to show that (20) preserves the symmetry properties of $dP$ and $\mathcal{E}$. Clearly, a symmetric $\mathcal{E}$ maps to a symmetric $dP$ by (19). What left is to show that (20) also maps a symmetric $dP$ to a symmetric $\mathcal{E}$. To begin with, let us introduce a permutation matrix $\Pi$ such that ${\rm vec}(A^{T})=\Pi\,{\rm vec}(A)$. Then, for any matrix $X$ with the same size as $P$, we have
$(P\otimes I+I\otimes P)\Pi\,{\rm vec}(X)=$ $(P\otimes I+I\otimes P){\rm vec}(X^{T})$
$=$ ${\rm vec}(X^{T}P+PX^{T})$
$=$ ${\rm vec}\left((PX+XP)^{T}\right)$
$=$ $\Pi{\rm vec}(PX+XP)$
$=$ $\Pi(I\otimes P+P\otimes I){\rm vec}(X)$
Since $X$ is arbitrary, the above equation suggests that
$\Pi^{-1}(P\otimes I+I\otimes P)=(I\otimes P+P\otimes I)\Pi^{-1}$ (21)
Then, from (20), we have
${\rm vec}(\mathcal{E}^{T})=$ $\Pi\,{\rm vec}(\mathcal{E})$
$=$ $\Pi(I\otimes P+P\otimes I)^{-1}{\rm vec}(dP)$
$=$ $\left((I\otimes P+P\otimes I)\Pi^{-1}\right)^{-1}{\rm vec}(dP)$
$=$ $\left(\Pi^{-1}(I\otimes P+P\otimes I)\right)^{-1}{\rm vec}(dP)$
$=$ $(I\otimes P+P\otimes I)^{-1}\Pi{\rm vec}(dP)$
$=$ $(I\otimes P+P\otimes I)^{-1}{\rm vec}(dP^{T})$
$=$ $(I\otimes P+P\otimes I)^{-1}{\rm vec}(dP)$
$=$ ${\rm vec}(\mathcal{E})$
where we have used (21) from the third to the fourth line. Hence, we have $\mathcal{E}=\mathcal{E}^{T}$ when $P\succ 0$. $\square$
By Proposition 3, (19) induces a new coordinate system for the manifold of SPD matrices. For a positive scalar $P$, integral on the right side of (20) gives the closed-form solution for transforming $P$ to the new one, i.e., $0.5\log P+{\rm constant}$. In general, we have no need to write down this transform explicitly for a matrix $P$, and (19) is enough for our purpose.
Now, we can rewrite the $dc(P;v,h)$ in (6) in the new coordinate as below after ignoring term $\mathcal{O}(\mathcal{E}^{3})$,
$\quad dc(P;v,h)$
$=h^{T}(P\mathcal{E}+\mathcal{E}P^{T})h-v^{T}P^{-1}(P\mathcal{E}+\mathcal{E}P^{T})P^{-1}v$
$\quad+v^{T}P^{-1}(P\mathcal{E}+\mathcal{E}P^{T})P^{-1}(P\mathcal{E}+\mathcal{E}P^{T})P^{-1}v$
$=h^{T}P\mathcal{E}h+h^{T}\mathcal{E}Ph-v^{T}\mathcal{E}P^{-1}v-v^{T}P^{-1}\mathcal{E}v+v^{T}\mathcal{E}^{2}P^{-1}v$
$\quad+v^{T}\mathcal{E}P^{-1}\mathcal{E}v+v^{T}P^{-1}\mathcal{E}P\mathcal{E}P^{-1}v+v^{T}P^{-1}\mathcal{E}^{2}v$ (22)
where in the last line we have assumed that $P=P^{T}$. Again, (22) is a quadratic function of $\mathcal{E}$, and we can solve for the $\mathcal{E}$ that reduces $dc(P;v,h)$ the most from the following equation,
$\quad Phh^{T}+hh^{T}P-vv^{T}P^{-1}-P^{-1}vv^{T}+vv^{T}P^{-1}\mathcal{E}^{T}$
$+\mathcal{E}^{T}vv^{T}P^{-1}+vv^{T}\mathcal{E}^{T}P^{-1}+P^{-1}\mathcal{E}^{T}vv^{T}$
$+P^{-1}vv^{T}P^{-1}\mathcal{E}^{T}P+P\mathcal{E}^{T}P^{-1}vv^{T}P^{-1}$
$+P^{-1}vv^{T}\mathcal{E}^{T}+\mathcal{E}^{T}P^{-1}vv^{T}=0$ (23)
Now, we can identify the gradient in the new coordinate from (23), and update $P$ with SGD given pair $(v_{t},h_{t})$ at step $t$ as,
$\mathcal{E}_{t}=$ $-\mu(P_{t}h_{t}h_{t}^{T}+h_{t}h_{t}^{T}P_{t}-v_{t}v_{t}^{T}P_{t}^{-1}-P_{t}^{-1}v_{t}v_{t}^{T})$
$P_{t+1}=$ $P_{t}+(P_{t}\mathcal{E}_{t}+\mathcal{E}_{t}P_{t})$ (24)
Clearly, with (24), $P_{t}$ always is symmetric as long as $P_{0}$ is.
Proposition 4: Starting from an initial guess $P_{0}\succ 0$ that commutes with $H$ and with a constant step size $0<\mu<0.5/\max_{t}{\lambda_{\max}(H+H^{2}P_{t})}$, the sequence $P_{t}$ given by (24) converges to $H^{-1}$ linearly with rate $1-8\mu\,\lambda_{\min}(H)$.
Proof: We take expectation on both sides of (24) to have
$\quad E[P_{t+1}]$
$=E[P_{t}]-\mu\left(E[P_{t}^{2}]H^{2}+2E[P_{t}H^{2}P_{t}]+H^{2}E[P_{t}^{2}]-4I\right)$
$=E[P_{t}]-\mu\left(\left(E[P_{t}]\right)^{2}H^{2}+2E[P_{t}]H^{2}E[P_{t}]\right.$
$\qquad\left.+H^{2}\left(E[P_{t}]\right)^{2}-4I+\mathcal{O}(\left(P_{t}-E[P_{t}]\right)^{2})\right)$ (25)
where $E[\cdot]$ is a shorthand for $E_{v_{1}\sim\mathcal{N}(0,I),\ldots,v_{t}\sim\mathcal{N}(0,I)}[\cdot]$. The term $\mathcal{O}(\left(P_{t}-E[P_{t}]\right)^{2})$ in the last line of (25) is negligible with a small $\mu$. We further drop the notation $E[\cdot]$ in (25), and just use $P_{t}$ to denote its expectation to simplify our writings.
First, we are to show that $P_{t+1}$ commutes with $H$ if $P_{t}$ does. Assuming $P_{t}H=HP_{t}$, we can simplify (25) to
$P_{t+1}=P_{t}-4\mu(H^{2}P_{t}^{2}-I)$ (26)
Then, it is not difficult to show that $P_{t+1}$ also commutes with $H$, similar to the proof of Proposition 2. Thus, $P_{t}$ always commutes with $H$ as long as $P_{0}$ does.
Second, let us define fitting error $R_{t}=HP_{t}-I$ and show that $\|R_{t}\|_{2}\rightarrow 0$ for $t\rightarrow\infty$. Clearly, $H$, $P_{t}$ and $R_{t}$ all are symmetric and commutes with each other. From (26), we have
$R_{t+1}=$ $HP_{t+1}-I$
$=$ $HP_{t}-4\mu H\left((HP_{t})^{2}-I\right)-I$
$=$ $(R_{t}+I)-4\mu H\left((R_{t}+I)^{2}-I\right)-I$
$=$ $(I-8\mu H-4\mu HR_{t})R_{t}$
$=$ $(I-4\mu H-4\mu H^{2}P_{t})R_{t}$ (27)
From the last line of (27), we see that $\mu$ needs to satisfy
$0<\mu<\frac{1}{2\lambda_{\max}(H+H^{2}P_{t})}$
to ensure the monotonic convergence of $P_{t}$. The second to the last line of (27) gives us the following asymptotic convergence rate
$1>\frac{\|R_{t+1}\|_{2}}{\|R_{t}\|_{2}}=\|I-8\mu H\|_{2}\geq 1-8\mu\,\lambda_{\rm min}(H)$
for $\|R_{t}\|_{2}\rightarrow 0$. Thus, $P_{t}$ converges linearly to $H^{-1}$ when $\lambda_{\rm min}(H)>0$. $\square$
Again, similar to the proof of Proposition 2, the condition that $P$ and $H$ commutes simplifies the proof, but could be too strict to be satisfied in practice. Yet, the linear convergence rate $1-8\mu\,\lambda_{\min}(H)$ matches well with observations.
Compared with SGD in the Euclidean space, linear convergence of the iterations given by (24) with a small enough constant step size is a good sign of progress. Intuitively, from (19), we see that $dP$ is roughly proportional to $P$ to mimic a Euclidean space varying step size adaptation. Following a similar way for the proof of Proposition 1, from (22), we can show that the Hessian in the new coordinates is given by
$S^{T}\left[3(P^{-1}\otimes I)+P^{-2}\otimes P\right]S$
which again is neither lower nor upper bounded. Thus, we have moved the needle a little bit, but not that much.
It is interesting to check the Newton’s methods in the new coordinates as well. Following the developments of the Newton’s methods in the Euclidean space, we can take expectation of the left side of (23) and solve for the optimal $\mathcal{E}$ directly. Generally, the two sets of Newton’s methods are different as the coordinate transform defined by (19) are not linear. However, if we assume that $\mathcal{E}$ and $P$ commutes, we can readily show that $dP$ and $P$ commutes as well, and eventually the two Newton’s methods are the same in this case.
II-E Hessian Fitting on Lie Groups
Linear convergence for Hessian fitting on Lie groups will not be a surprise as a connected Lie group is a special manifold equipped with extra properties that can further facilitate the development of Hessian fitting methods. One significant advantage of fitting $P$ on certain Lie groups is that they enable us to tightly lower bound the Hessian of $c(P)$. We leave these developments to the next section.
II-F Summary of All the Hessian Fitting Methods
We summarize most of the above candidates for Hessian fitting in Table 2. To confirm our analyses, Fig. 1 shows a set of typical convergence curves when $H$ is a small Hilbert matrix. Except for the Newton method that starts from $P_{0}=0.02I$, all the other methods start from $P_{0}=I$. The step sizes of SGD in the Euclidean space and the manifold of SPD matrices are tuned to achieve the maximum rate of convergence. Step sizes for SGD on Lie group and Newton’s method are already normalized to $1$. Except for the SGD (10) in the Euclidean space and the closed-form solution (17) that converge sublinearly, all the other methods converge to $H^{-1}$ within reasonable numbers of iterations. These empirical convergence rates match well with the predictions of Table 2.
TABLE II: Candidates for Hessian fitting. Note that $10<\rho<$ 1 is not necessarily the same for different methods.
Method Equation Space Error at $t$
SGD (10) Euclidean $\mathcal{O}(1/\sqrt{t})$
Closed-form (17) Euclidean $\mathcal{O}(1/t)$
SGD (24) SPD $\mathcal{O}(\rho^{t})$
SGD (36) Lie group $\mathcal{O}(\rho^{t})$
Newton (14) Euclidean/SPD $\mathcal{O}(\rho^{2^{t}})$
Figure 1: (a): Typical convergence curves for the five Hessian fitting methods in Table 2 when $H$ is a $3\times 3$ Hilbert matrix, i.e., $H_{i,j}=1/(i+j-1)$ with $1\leq i,j\leq n$. (b) and (c): Zooming in of the converge curves of SGD on Lie group and Newton’s method for better visualization, respectively.
III Hessian Fitting on Lie Groups
III-A Strong Convexity on Certain Lie Groups
It is still not handy to fit Hessian on the manifold of SPD matrices as we need to constantly guard the positive definiteness of $P_{t}$. Also, per Proposition 4, knowledge of $H$ is required to determine the fitting step sizes. Here, we let
$P=Q^{T}Q$ (28)
[…]
With the Lie group constraint, we let $I+\mathcal{E}=e^{\varepsilon G}$, where $\varepsilon\rightarrow 0$ is a small scalar, and $G$ is the group generator. Then, we can have the following two choices for $dQ$,
${\rm choice\,I:}\quad$ $dQ=e^{\varepsilon G}Q-Q=\mathcal{E}Q$
${\rm choice\,II:}\quad$ $dQ=Qe^{\varepsilon G}-Q=Q\mathcal{E}$ (29)
Correspondingly, the $dP$ in (28) will have the following two possible forms,
${\rm choice\,I:}\quad dP=$ $(Q+\mathcal{E}Q)^{T}(Q+\mathcal{E}Q)-P$
$=$ $Q^{T}(\mathcal{E}+\mathcal{E}^{T}+\mathcal{E}^{T}\mathcal{E})Q$
${\rm choice\,II:}\quad dP=$ $(Q+Q\mathcal{E})^{T}(Q+Q\mathcal{E})-P$
$=$ $\mathcal{E}^{T}P+P\mathcal{E}+\mathcal{E}^{T}P\mathcal{E}$ (30)
The second choice of $dP$ in (30) looks almost the same as that in (19) after ignoring the second order term of $\mathcal{E}$. But, now $\mathcal{E}$ is a small matrix proportional to the group generator, and no longer required to be symmetric. Substituting (30) back to (6) and noting that $P^{-1}=Q^{-1}Q^{-T}$, we can arrive at the following two possible choices for $dc(P;v,h)$,
$dc_{I}(P;v,h)=h^{T}Q^{T}(\mathcal{E}+\mathcal{E}^{T})Qh-v^{T}Q^{-1}(\mathcal{E}+\mathcal{E}^{T})Q^{-T}v$
$\qquad\quad+h^{T}Q^{T}\mathcal{E}^{T}\mathcal{E}Qh+v^{T}Q^{-1}(\mathcal{E}^{2}+\mathcal{E}\mathcal{E}^{T}+\mathcal{E}^{2T})Q^{-T}v$
$dc_{II}(P;v,h)=h^{T}(\mathcal{E}^{T}P+P\mathcal{E})h-v^{T}(P^{-1}\mathcal{E}^{T}+\mathcal{E}P^{-1})v$
$\qquad\quad+h^{T}\mathcal{E}^{T}P\mathcal{E}h+v^{T}(\mathcal{E}^{2}P^{-1}+\mathcal{E}P^{-1}\mathcal{E}^{T}+P^{-1}\mathcal{E}^{2T})v$ (31)
where we have suppressed higher order term $\mathcal{O}(\mathcal{E}^{3})$ to shorten our equations. The first choice in (31) not only looks more balanced and symmetric, but also have nicer properties than the second one as will be revealed in our following discussions. Thus, we mainly focus on the first choice.
One striking difference between fitting Hessian on the Lie group and the manifold of SPD matrices is that criterion $c(P)$ could be strongly convex on a few Lie groups under certain mild conditions like $\lambda_{\min}(H)>0$. Clearly, $c(P)$ is not convex on GL$(n,\mathbb{R})$ as $Q$ and $UQ$ give the same value for $c(P)$ for any orthogonal matrix $U$. But, $c(P)$ could behave better after properly separating out this rotation ambiguity. To this end, we need to turn to the polar decomposition [24] $A=U\Gamma$ to define a proper Lie group that is irrelevant of such rotations, where $A\in\mathbb{R}^{n\times n}$ is a square matrix, $U$ is an orthogonal matrix, and $\Gamma\succeq 0$ is semi-positive definite. With the polar decomposition, let us define the following equivalence relation.
Equivalence relation $R_{\rm polar}$: For $Q_{1}$ and $Q_{2}$ in $\mathbb{R}^{n\times n}$, we say that $Q_{1}\sim Q_{2}$ if and only if their polar decompositions share the same $\Gamma$, i.e., $Q_{1}=U_{1}\Gamma$ and $Q_{2}=U_{2}\Gamma$ for certain orthogonal matrices $U_{1}$ and $U_{2}$.
Then, we can define the quotient set GL$(n,\mathbb{R})/R_{\rm polar}$. It is not difficult to show that all the elements of GL$(n,\mathbb{R})/R_{\rm polar}$ form a Lie group. We skip the proof here due to its straightforwardness. Note that since group $O(n)$ has dimension $0.5n(n-1)$, elements from this group have dimension $n^{2}-0.5n(n-1)=0.5n(n+1)$. However, we cannot represent them as SPD matrices since set $\{A|A\succ 0\}$ is not closed with matrix multiplication operation. Abstract mathematical properties of this group are not interested here.
Proposition 5: $c(P)$ is strongly convex on group GL$(n,\mathbb{R})/R_{\rm polar}$ with the coordinate system induced by choice I in (29) when $\lambda_{\min}(H)\geq\lambda_{0}>0$.
Proof: First, let us briefly show that with $A\succ 0$ and eigenvalue decomposition $A=UDU^{T}$, $aA+bA^{-1}-2\sqrt{ab}I=U(aD+bD^{-1}-2\sqrt{ab})U^{T}\succeq 0$, where $a$ and $b$ are two positive numbers.
Second, we show that it is sufficient to consider a symmetric $\mathcal{E}$ on group GL$(n,\mathbb{R})/R_{\rm polar}$. Starting from $Q$, let $Q_{1}=Q+\mathcal{E}_{1}Q$ be a neighbor of $Q$ on GL$(n,\mathbb{R})/R_{\rm polar}$ with an arbitrary $\mathcal{E}_{1}$ small enough such that $\|\mathcal{E}_{1}\|_{2}<1$. Let the polar decomposition of $Q_{1}$ and $I+\mathcal{E}_{1}$ be
$Q_{1}=U_{1}\Gamma_{1},\qquad I+\mathcal{E}_{1}=U_{2}\Gamma_{2}$
respectively. Then, we have
$\Gamma_{2}Q=U_{2}^{T}(I+\mathcal{E}_{1})Q=U_{2}^{T}Q_{1}=(U_{2}^{T}U_{1})\Gamma_{1}$
Thus, by letting $\mathcal{E}_{2}=\Gamma_{2}-I$, we have
$(I+\mathcal{E}_{2})Q=(U_{2}^{T}U_{1})\Gamma_{1}\sim U_{1}\Gamma_{1}=(I+\mathcal{E}_{1})Q$ (32)
Since $\Gamma_{2}$ is symmetric, $\mathcal{E}_{2}=\Gamma_{2}-I$ is symmetric as well. Thus, (32) suggests that it is enough to only consider a symmetric $\mathcal{E}$ on group GL$(n,\mathbb{R})/R_{\rm polar}$. This is not a surprise since a symmetric $\mathcal{E}$ just has dimension $0.5n(n+1)$.
The last step is to check the lower bound of the second order terms of $dc_{I}(P)$ in (31). Now, without loss of generality, let us assume that $\mathcal{E}=\mathcal{E}^{T}$ due to (32). Starting from (31), the expectation of the second order terms of $dc_{I}(P;v,h)$ with choice $dQ=\mathcal{E}Q$ is given by
$E_{v}\left[h^{T}Q^{T}\mathcal{E}^{T}\mathcal{E}Qh+v^{T}Q^{-1}(\mathcal{E}^{2}+\mathcal{E}\mathcal{E}^{T}+\mathcal{E}^{2T})Q^{-T}v\right]$
$=$ ${\rm tr}\left(\mathcal{E}^{2}QH^{2}Q^{T}+3\mathcal{E}^{2}Q^{-T}Q^{-1}\right)$
$\geq$ ${\rm tr}\left(\left(\lambda_{\min}(H)\right)^{2}\mathcal{E}^{2}QQ^{T}+3\mathcal{E}^{2}\left(QQ^{T}\right)^{-1}\right)$
$\geq$ ${\rm tr}\left(\mathcal{E}^{2}\left(\lambda_{0}^{2}QQ^{T}+3(QQ^{T})^{-1}\right)\right)$
$\geq$ $2\sqrt{3}\lambda_{0}\,{\rm tr}\left(\mathcal{E}^{2}\right)$
$=$ $2\sqrt{3}\lambda_{0}\left({\rm vec}(\mathcal{E})\right)^{T}{\rm vec}(\mathcal{E})$
$=$ $2\sqrt{3}\lambda_{0}\left({\rm vech}(\mathcal{E})\right)^{T}S^{T}S\,{\rm vech}(\mathcal{E})$
$=$ $\left({\rm vech}(\mathcal{E})\right)^{T}\left(2\sqrt{3}\lambda_{0}\,S^{T}S\right){\rm vech}(\mathcal{E})$
where we have used expectations $E_{v}[hh^{T}]=H^{2}$ and $E_{v}[vv^{T}]=I$ from the first to the second line, $aA+bA^{-1}\succ 2\sqrt{ab}I$ for $A\succ 0$ to arrive the fifth line from the fourth one, and definition ${\rm vec}(A)=S{\rm vech}(A)$ for a symmetric matrix $A$ to arrive the last line. Thus, the Hessian of $c_{I}(P)$ in the new coordinate is lower bounded by $2\sqrt{3}\lambda_{0}\,S^{T}S\succ 2\sqrt{3}\lambda_{0}I$ since $S^{T}S\succ I$. $\square$
We briefly present a few more statements regarding the convexity of Hessian fitting on Lie groups without formal proofs since they are less important. Clearly, $c(P)$ is strongly convex on any connected group of diagonal matrices when $\lambda_{\min}(H)\geq\lambda_{0}>0$. On group $O(n)$, from (31), it is not difficult to show that every point is a saddle point as $\mathcal{E}=-\mathcal{E}^{T}$. The second choice, $c_{II}(P)$, is generally not convex even with constraint $\mathcal{E}=\mathcal{E}^{T}$. Empirical results also have confirmed that Hessian fitting with choice II converges significantly slower than that with choice I. The first choice, $c_{I}(P)$, is generally not convex on the group of triangular matrices, although previous results have shown that Hessian fitting on this group does work very well [17]. An explanation is given in the next subsection.
III-B The Two Baseline Lie Group Hessian Fitting Methods
III-B1 Fitting on Group GL$(n,\mathbb{R})/R_{\rm polar}$
As our Hessian fitting criterion is strongly convex on this group, we can readily follow the conventional convex optimization theory to design proper Hessian fitting methods [1]. If we could find an upper bound for the eigenvalues of the Hessian of $c(P)$, say $L$, we know that gradient descent with step size $1/L$ converges linearly to the optimal solution $H^{-1}$ at least with rate $1-2\sqrt{3}\lambda_{0}/L$, where $2\sqrt{3}\lambda_{0}$ is the lower bound of Hessian derived in Proposition 5. From (31), choice I, we know that sum of the second order terms of $\mathcal{E}$ is upper bounded by
$\left|h^{T}Q^{T}\mathcal{E}^{T}\mathcal{E}Qh+v^{T}Q^{-1}(\mathcal{E}^{2}+\mathcal{E}\mathcal{E}^{T}+\mathcal{E}^{2T})Q^{-T}v\right|$
$\leq$ $\left|h^{T}Q^{T}\mathcal{E}^{T}\mathcal{E}Qh\right|+\left|v^{T}Q^{-1}\mathcal{E}^{2}Q^{-T}v\right|$
$+\left|v^{T}Q^{-1}\mathcal{E}\mathcal{E}^{T}Q^{-T}v\right|+\left|v^{T}Q^{-1}\mathcal{E}^{2T}Q^{-T}v\right|$
$\leq$ $\left(\|Qh\|^{2}+3\|Q^{-T}v\|^{2}\right)\|\mathcal{E}\|^{2}$ (33)
where $\|\mathcal{E}\|$ can be any matrix norm for $\mathcal{E}$. Thus, from (33), we see that an extremely rough estimate of $L$ could be
$L_{t}\approx\max_{1\leq i\leq t}\left(\|Q_{i}h_{i}\|^{2}+3\|Q_{i}^{-T}v_{i}\|^{2}\right)$ (34)
At the same time, from the linear terms in (31), choice I, we see that the gradient on group GL$(n,\mathbb{R})$ is
$\nabla_{Q}=2\left(Qhh^{T}Q^{T}-Q^{-T}vv^{T}Q^{-1}\right)$ (35)
which also is the gradient on group GL$(n,\mathbb{R})/R_{\rm polar}$ since it is already symmetric. Combining (33) and (35) gives the following first baseline Lie group SGD Hessian fitting method,
$Q_{t+1}=Q_{t}-\mu\frac{Q_{t}h_{t}h_{t}^{T}Q_{t}^{T}-Q_{t}^{-T}v_{t}v_{t}^{T}Q_{t}^{-1}}{\|Q_{t}h_{t}\|^{2}+\|Q_{t}^{-T}v_{t}\|^{2}}Q_{t}$ (36)
where $0<\mu\leq 1$ is a normalized step size. Note that in (36), we do not use the more conservative $L$ estimate given by (34). Numerical results suggest that (36) is stable enough.
A note on the matrix inverse in (36). Straightforward implementation of (36) needs to solve the linear system $Q^{T}x=v$ for calculating $Q^{-T}v$. Still, from (36), we see that $Q_{t+1}-Q_{t}$ is a matrix with rank two. Thus, we can recursively update the inverse of $Q$ with the Woodbury matrix identity [24] as in BFGS. A concern is that the numerical errors may accumulate after repeatedly applying the Woodbury matrix identity many times. Empirical results have confirmed that this procedure is generally numerically stable with double and single precision arithmetic, but not stable with half precision arithmetic.
III-B2 Fitting on the Group of Triangular Matrices
We can simply take the triangular part of (35) to get the gradient for updating $Q$ on the group of triangular matrices as having done in [17]. But, in general $c(P)$ is not convex on this group. To take the advantage of Proposition 5, we can factorize out the rotation part of the updated $Q$ from (36) to have the following update formula for $Q$ on the group of upper triangular matrix,
$Q_{t+1}=\left[I-\mu\frac{Q_{t}h_{t}h_{t}^{T}Q_{t}^{T}-Q_{t}^{-T}v_{t}v_{t}^{T}Q_{t}^{-1}}{\|Q_{t}h_{t}\|^{2}+\|Q_{t}^{-T}v_{t}\|^{2}}\right]_{R}Q_{t}$ (37)
where $[A]_{R}$ only keeps the triangular part of the QR decomposition of a square matrix $A$. In this way, $Q_{t}$ is still on the group GL$(n,\mathbb{R})/R_{\rm polar}$, but keeps to be a triangular matrix as long as $Q_{0}$ is. Note that only two rank-1 QR decomposition updates are required to obtain the QR decomposition of the matrix inside $[\cdot]_{R}$ in (37) due to its low-rank structure [24]. One pleasant by-product here is that now, we can calculate $Q_{t}^{-T}v_{t}$ with backward substitution, which is empirically shown to be numerically stable even with half precision arithmetic.
Let us derive an approximation for $[A]_{R}$ to reveal the connection between (37) and the method in [17]. Let $\|\Delta\|<1$ be a small matrix, and consider the following QR decomposition,
$I+\Delta$ $=(I+\Delta_{1})(I+\Delta_{2})$
$I$ $=(I+\Delta_{1})^{T}(I+\Delta_{1})$ (38)
where $I+\Delta_{1}$ is an orthogonal matrix, and $\Delta_{2}$ is an upper triangular matrix. Note that $\Delta_{1}$ and $\Delta_{2}$ are small matrices as well since QR decomposition is numerically stable. After expanding (38) and ignoring the second order terms $\Delta_{1}^{T}\Delta_{1}$ and $\Delta_{1}\Delta_{2}$, we can solve for $\Delta_{1}$ and $\Delta_{2}$ jointly from equations,
$\Delta_{1}+\Delta_{2}\approx\Delta,\qquad\Delta_{1}+\Delta_{1}^{T}\approx 0$
which can be shown to be,
$\Delta_{1}$ $\approx{\rm tril}(\Delta,-1)-{\rm triu}(\Delta,1)$
$\Delta_{2}$ $\approx{\rm triu}(\Delta)+{\rm triu}(\Delta,1)$
where ${\rm tril}(A,-1)$ and ${\rm triu}(A,1)$ exclude the diagonal elements of ${\rm tril}(A)$ and ${\rm triu}(A)$, respectively. Hence, we obtain the following approximation for $[A]_{R}$,
$\left[I+\Delta\right]_{R}\approx{\rm triu}(\Delta)+{\rm triu}(\Delta,1),\qquad{\rm for}\,\|\Delta\|<1$ (39)
In this way, we see that gradient on the group of triangular matrices still points to an ascent direction on the group GL$(n,\mathbb{R})/R_{\rm polar}$, although not the steepest one. We always take approximation (39) for the $[\cdot]_{R}$ operation in (37). Numerical results have confirmed that after separating out this rotation ambiguity, update (37) roughly converges twice faster than the one given in [17].
III-B3 Newton’s Methods on Lie Groups
Lastly, we want to briefly mention the Newton’s methods on Lie groups, which can be derived from (31) by solving for the optimal $\mathcal{E}$ that reduces $c(P)$ the most. But, they turn out to be way more complicated than the form in (14), and seem to be of little practical value. Nevertheless, the Hessian for the case considered in Proposition 5 still has the following neat form,
$S^{T}\left[I\otimes(QH^{2}Q^{T}+3Q^{-T}Q^{-1})\right]S$
III-C Sparse Lie Groups for Large Scale Hessian Fittings
The methods in (36) and (37) are only good for small to middle scale problems with roughly $n\leq 10^{5}$ parameters to optimize. For problems with millions to billions of parameters to learn, we need significantly sparser Hessian estimators.
III-C1 The Group of Diagonal Matrices
Let $Q={\rm diag}(q)$, where $q$ is a vector without zero element. Note that $\|A\|_{2}=\max_{1\leq i\leq n}|a_{i,i}|$ when $A$ is diagonal. It is trivial to following the same procedures leading to (36) to derive the following SGD method for updating $Q$ on this simple group,
$q_{t+1}=q_{t}-\mu\frac{\left(q_{t}h_{t}\right)^{2}-\left(v_{t}/q_{t}\right)^{2}}{\max\left(\left(q_{t}h_{t}\right)^{2}+\left(v_{t}/q_{t}\right)^{2}\right)}q_{t}$ (40)
III-C2 Decomposable Groups
A $Q$ that can be decomposed into form $Q=\oplus_{i=1}^{I}\left(\otimes_{j=1}^{J_{i}}Q_{i,j}\right)$ can be useful for neural network trainings since those models mainly consist of a list of modules using matrices or tensors as trainable parameters, see [18] for a detailed treatment. Specifically, let $\Theta$ be a matrix parameter, and $y=\Theta x$ be the output of a certain layer with input feature vector $x$, where the last element $x$ is $1$ such that this is a general affine transform. Then, the PSGD iteration in (1) with $Q=Q_{2}\otimes Q_{1}$ equals an SGD update in a $\Theta$-dependent new coordinate defined by $\Theta^{\prime}=Q_{1}^{-T}\Theta Q_{2}^{-1}$ with affine transform $y^{\prime}=\Theta^{\prime}x^{\prime}$, where $y^{\prime}=Q_{1}^{-T}y$ and $x^{\prime}=Q_{2}x$ [18]. Hence, we call this type of preconditioner the affine one. Let us derive an improved implementation for fitting such preconditioners in the light of Proposition 5.
Without loss of generality, we only consider the case of $Q=Q_{2}\otimes Q_{1}$. First, let us freeze $Q_{1}$ and only consider the update of $Q_{2}$. By Proposition 5, we should take $dQ_{2}=\mathcal{E}_{2}Q_{2}$, and thus,
$dQ=\left(\mathcal{E}_{2}Q_{2}\right)\otimes Q_{1}=\left(\mathcal{E}_{2}\otimes I\right)\left(Q_{2}\otimes Q_{1}\right)=\left(\mathcal{E}_{2}\otimes I\right)Q$
Then, by replace the $\mathcal{E}$ in (31) with $\left(\mathcal{E}_{2}\otimes I\right)$, we obtain the second order approximation for $c(P+dP)-c(P)$. Also, noting that $\|\mathcal{E}_{2}\otimes I\|_{2}=\|\mathcal{E}_{2}\|_{2}$, we can follow the procedures leading to (37) to bound the Hessian with respect to $\mathcal{E}_{2}$ to derive a step size normalized update rule for $Q_{2}$. In the same way, we can freeze $Q_{2}$ and have $dQ=\left(I\otimes\mathcal{E}_{1}\right)Q$ by taking $dQ_{1}=\mathcal{E}_{1}Q_{1}$. Again, noting that $\|I\otimes\mathcal{E}_{1}\|_{2}=\|\mathcal{E}_{1}\|_{2}$, we can derive another step size normalized update rule for $Q_{1}$. We skip these tedious details, and directly give the SGD update for $Q$ as below,
$A$ $=Q_{1,t}{\rm uvec}(h_{t})Q_{2,t}^{T},\;B=Q_{2,t}^{-T}\left[{\rm uvec}(v_{t})\right]^{T}Q_{1,t}^{-1}$
$Q_{1,t+1}$ $=\left[I-\mu\,\frac{AA^{T}-B^{T}B}{\|A\|_{F}^{2}+\|B\|_{F}^{2}}\right]_{R}Q_{1,t}$
$Q_{2,t+1}$ $=\left[I-\mu\,\frac{A^{T}A-BB^{T}}{\|A\|_{F}^{2}+\|B\|_{F}^{2}}\right]_{R}Q_{2,t}$ (41)
where ${\rm uvec}(v_{t})$ and ${\rm uvec}(h_{t})$ recover vectors $v_{t}$ and $h_{t}$ to their original matrix forms, respectively.
Eq. (41) looks almost the same as the one in [17] in Section V.B for updating $Q$ except that here, we normalize the gradient with $\|A\|_{F}^{2}+\|B\|_{F}^{2}$ and remove the rotation ambiguity by using the operator $[\cdot]_{R}$ approximation in (39). Again, in practice, the matrix inverses in (41) should be replaced with backward substitutions, which is as cheap and numerically stable as matrix multiplications. Also note that (41) reduces to (37) by choosing $Q_{2}=1$ when $\Theta$ is a column vector.
Note that the KFAC [5] and Shampoo [6, 7, 8, 9] like preconditioners are fundamentally different from ours even though they all share the same Kronecker product forms. The KFAC and Shampoo preconditioners can only approximate the inverse of Hessians even when they can be factorized into the assumed Kronecker product forms, while our affine one could converge to the optimal solution given the correct factorization form.
III-C3 Low-Rank Approximation (LRA) Preconditioner
Similar to the L-BFGS method, we have developed a LRA preconditioner in [19]. Let us revisit it and refine its implementations. We can have two choices for the definition of $Q$,
${\rm choice\,I:}\quad Q$ $=(I+UV^{T}){\rm diag}(d)$
${\rm choice\,II:}\quad Q$ $={\rm diag}(d)(I+UV^{T})$ (42)
where vector $d$ has no zeros such that ${\rm diag}(d)$ is an element of the group of diagonal matrices, and $U$ and $V\in\mathbb{R}^{n\times r}$ with $0\leq r\ll n$. Both choices have the same representation capacity for nonsingular matrices with form ${\rm diag}(d^{\prime})+U^{\prime}V^{\prime T}$. Still, the first choice is easier to work with when we are to bound the Hessian with respect to $U$ and $V$.
The key to work with the LRA preconditioner is to notice the Lie group structure of the $Q$ in (42). Clearly, ${\rm diag}(d)$ is from the group of diagonal matrices. Noting that
$e^{\varepsilon G_{1}V^{T}}=I+\left(\sum_{i\geq 1}\frac{\varepsilon^{i}(G_{1}V^{T})^{i-1}G_{1}}{i!}\right)V^{T}=\mathcal{E}_{1}V^{T}$ (43)
we see that set $\{I+UV^{T}|U\in\mathbb{R}^{n\times r},V{\rm\,is\,fixed}\}$ forms a Lie group. Similarly, with a given $U$, set $\{I+UV^{T}|V\in\mathbb{R}^{n\times r},U{\rm\,is\,fixed}\}$ forms another Lie group. We can exploit these Lie group structures to update $U$, $V$ and $d$ efficiently. For example, to update $U$, from (43), we can let $dQ$ be
$dQ=e^{\varepsilon G_{1}V^{T}}Q-Q=\mathcal{E}_{1}V^{T}Q$
and then simply replace the $\mathcal{E}$ in (31) with $\mathcal{E}_{1}V^{T}$ to obtain the second order approximation of $c(P+dP)-c(P)$. Similarly, to update $V$, we switch to the other group and let $dQ$ be
$dQ=e^{\varepsilon UG_{2}^{T}}Q-Q=U\mathcal{E}_{2}^{T}Q$
We omit the lengthy derivations here and only present the final update rules for $U$, $V$ and $d$ as below,
$a=$ $Q_{t}h_{t},\quad b=Q_{t}^{-T}v_{t}$
$d_{t+1}=$ $\left(1-\mu\frac{(Q_{t}^{T}a)h_{t}-v_{t}(Q_{t}^{-1}b)}{\max\sqrt{\left((Q_{t}^{T}a)^{2}+v_{t}^{2}\right)\left(h_{t}^{2}+(Q_{t}^{-1}b)^{2}\right)}}\right)d_{t}$
$U_{t+1}=$ $U_{t}-\mu\frac{(aa^{T}-bb^{T})V_{t}(I+V_{t}^{T}U_{t})}{\|a\|\|V_{t}V_{t}^{T}a\|+\|b\|\|V_{t}V_{t}^{T}b\|}$
$V_{t+1}=$ $V_{t}-\mu\frac{(I+V_{t}U_{t}^{T})(aa^{T}-bb^{T})U_{t}}{\|a\|\|U_{t}U_{t}^{T}a\|+\|b\|\|U_{t}U_{t}^{T}b\|}$ (44)
Interested readers can refer to the supplementary materials of [19] for further details. The only difference is that in (44), we have normalized the step sizes with second order derivatives.
A few notes before ending the LRA preconditioner discussions. The matrix inverse in (44) can be cheaply calculated with the Woodbury matrix identity [24] since typically $0\leq r\ll n$. We usually update either $U$ and $V$, not both, within a single step since set $\{I+UV^{T}|U,V\in\mathbb{R}^{n\times r}\}$ no longer forms a Lie group. Our LRA preconditioner is quite different from the L-BFGS method. The L-BFGS is basically a limited number steps of expansion of the BFGS iterations, and has no long term memories. Clearly, (44) is not a simple expansion of (36) up to $r$ steps. It keeps tracking of the best approximation of $H^{-1}$ with a diagonal plus a low-rank matrices. Also, it is not difficult to check that with $Q={\rm diag}(q)$,
$\left((Q_{t}^{T}a)^{2}+v_{t}^{2}\right)\left(h_{t}^{2}+(Q_{t}^{-1}b)^{2}\right)=\left(\left(q_{t}h_{t}\right)^{2}+\left(v_{t}/q_{t}\right)^{2}\right)^{2}$
which suggests that (44) reduces to (40) when $r=0$.
III-C4 More Sparse Lie Groups for Preconditioner Fittings
It is not possible to enumerate all the potentially useful sparse Lie groups for preconditioner fittings. One systematic way to create such sparse Lie groups is to construct the so-called matrix-free preconditioners from any subgroups of the (discrete) permutation group with order $n$ [19]. Specifically, the diagonal preconditioner is the simplest case from this family, induced by the identity element, $\{e\}$, from the permutation group. For an even $n$, $\{e,s_{n/2}\}$ induces the butterfly matrices which also form a Lie group, where $s_{n/2}$ denotes the $n/2$-circular-shift operator [19]. All these groups can be reduces to the direct sum of a series of small ones.
Another interesting and nontrivial Lie group is the set of matrices having form [18]
$Q=\begin{bmatrix}{\rm bdiag}(A_{1},\ldots,A_{i})&B\\ 0&C\end{bmatrix}$ (45)
where ${\rm bdiag}(A_{1},\ldots,A_{i})$ is a block diagonal matrix, $C$ and all the blocks $A_{j}$, $1\leq j\leq i$, are invertible, and $B$ is a matrix with proper size. It is called the normalization Lie group in [18] since as a special affine transform, a feature normalization matrix does have this form. For example, in the batch [15] or layer normalizations [16], a feature vector $x$ is partitioned into $i$ groups as $x=[x_{1};\ldots;x_{i};1]$ and normalized as,
$\begin{bmatrix}x_{1}^{\prime}\\ \vdots\\ x_{i}^{\prime}\\ 1\end{bmatrix}=\begin{bmatrix}\frac{1}{a_{1}}I&&&-\frac{b_{1}}{a_{1}}e\\ &\ddots&&\vdots\\ &&\frac{1}{a_{i}}I&-\frac{b_{i}}{a_{i}}e\\ &&&1\end{bmatrix}\begin{bmatrix}x_{1}\\ \vdots\\ x_{i}\\ 1\end{bmatrix}$ (46)
where $a_{j}$ and $b_{j}$, $1\leq j\leq i$, are the shared variance and mean of the $j$th group, respectively, and $e$ is a vector of ones. It is not difficult to check that matrices with the form in (46) do form a Lie group. Due to the importance of feature normalizations in neural network learning [15, 16], this simple Lie group is useful by providing a tool for fitting such preconditioners in an elegant and principled way.
III-D Summary of Lie Group Hessian Fitting Methods
Table III summarizes the information of the above five basic types of Lie group preconditioners. With the diagonal preconditioner as the base line and assuming that $\theta={\rm vec}(\Theta)$ with $\Theta\in\mathbb{R}^{m\times m}$, we see that the affine preconditioner, i.e., $Q=Q_{2}\otimes Q_{1}$, requires more computations, and the LRA one with $r\ll m$, i.e., $Q=(I+UV^{T}){\rm diag}(d)$, needs more memory space. It is possible to mix these basic building blocks to derive richer forms of Lie group preconditioners, e.g., the scaling-normalization preconditioner in [18] which only has a sublinear complexity, i.e., $\mathcal{O}(m)$.
TABLE III: Hessian fitting on Lie groups. The storages and computations are the numbers assuming that $\theta={\rm vec}(\Theta)$ with $\Theta\in\mathbb{R}^{m\times m}$.
$Q$ $Q_{t+1}\leftarrow Q_{t}$ Storages Computations
GL$(n,\mathbb{R})$ (36) $\mathcal{O}(m^{4})$ $\mathcal{O}(m^{4})$
Triangular (37) $\mathcal{O}(m^{4})$ $\mathcal{O}(m^{6})$
Diagonal (40) $\mathcal{O}(m^{2})$ $\mathcal{O}(m^{2})$
$Q_{2}\otimes Q_{1}$ (41) $\mathcal{O}(m^{2})$ $\mathcal{O}(m^{3})$
$(I+UV^{T}){\rm diag}(d)$ (44) $\mathcal{O}(rm^{2})$ $\mathcal{O}(rm^{2})$
Fig. 2 shows a few typical Hessian fitting convergence curves for the PSGD methods with (36), (37) and default step size $\mu=1$, the closed-form solution (17) by clipping the maximum moving average factor to $0.999$, and the classic BFGS formula [1]. We have tested the following three cases.
III-D1 Clean Hessian-Vector Products
In this case, we set the ground truth $H^{\prime}$ to $H$. All methods except for the closed-form solution can converge to solutions with errors proportional to the machine precision. The PSGD style with (36) and BFGS have higher error plateaus than the PSGD style with (37) due to the accumulated numerical errors after repeated usage of the Woodbury matrix identity [24] for recursive matrix inverse.
III-D2 Noisy Hessian-Vector Products
We have tested a noisy model of (3) with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}I)$, and $\sigma_{\epsilon}=0.01$. Hence, we set the ground truth to $H^{\prime}=(H^{2}+\sigma_{\epsilon}^{2}I)^{0.5}$ due to (5). We see that the BFGS is very sensitive to noises, and diverges easily. The two PSGD methods with reduced step size $\mu=0.1$ converge slower than the closed-form solution, but reach lower fitting errors. Note that now there always are steady state misadjustment errors with PSGD since the gradient in (35) cannot settle down to zero when $\sigma_{\epsilon}>0$. Yet, Proposition 5 still holds after replacing the $H$ there with $=(H^{2}+\sigma_{\epsilon}^{2}I)^{0.5}$. The fitting errors of PSGD can be sufficiently small given enough iterations to anneal the step sizes.
III-D3 Time-Varying Hessians
The Hessians are from a time varying process defined by $H_{t+1}=H_{t}+uu^{T}$, where $u_{i}\sim\mathcal{U}(0,1)$ for $1\leq i\leq n$, and $H_{0}$ is a matrix with all elements being $1/4$. The ground truth $H^{\prime}=H_{t}$ is also time-varying. The BFGS diverges first as the Hessians changes too fast, and only locks onto the target $H_{t}^{-1}$ after $t>1000$. The closed-form solution has difficulty in tracking a time-varying Hessian due to its sublinear convergence. The two PSGD style Hessian fitting methods lock onto the target quicker than BFGS, and also do not show a divergence phase before convergence.
Figure 2: (a) Typical convergence curves when $H$ is a $50\times 50$ matrix with $H_{i,i}=1$ and $H_{i,j}=0.5$ for $|i-j|=1$. Its eigenvalues are bounded in range $(0,2)$. (b) The same $H$ as in (a), but with a noisy model of (3), where $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}I)$, and $\sigma_{\epsilon}=0.01$. (c) Time varying Hessians defined by process $H_{t+1}=H_{t}+uu^{T}$, where $u_{i}\sim\mathcal{U}(0,1)$ for $501\leq i\leq n=$ 50, and $H_{0}$ is a matrix with all elements being $1/4$.
IV More Experimental Results
So far we have detailed the behaviors of a wide range of Hessian fitting methods, see the summaries in Table II and III, and Fig. 1 and Fig. 2. Here, we present a few more experiments results to demonstrate their typical behaviors in practice. Interested readers can refer to our reports in [21] for significantly more extensive comparison results among these Hessian fitting methods and many more optimizers on a wide range of machine learning benchmarks. Focus here is on the theoretical properties of different stochastic Hessian fitting methods. In general, we observe that accurate Hessian estimation is strongly correlated with faster convergences and better generalizations. In this section, the SGD only refers to the one for learning the parameter $\theta$, not the one for Hessian fitting. The PSGD refers to the improved versions in this paper, not the originals ones from [17, 18, 19]. For PSGD, the default step size for preconditioner fitting is $0.1$.
IV-A PSGD vs. Approximated Closed-Form Solution Shampoo
This example compares PSGD with the Shampoo [7] on the LeNet5 [25] convolutional network for handwriting digits recognitions. With gradient $G={\rm uvec}(g)$ for matrix parameter $\Theta$, Shampoo uses closed-form solution
$\left(E[G^{T}G]\right)^{-0.25}\otimes\left(E[GG^{T}]\right)^{-0.25}$ (47)
as the preconditioner. Note that (47) is not an exact solution even when $E[gg^{T}]$ can be exactly factorized into a Kronecker product of two matrices. Fig. 3 shows the typical convergence curves of SGD, Shampoo and PSGD-Affine with batch size $64$ and proper tunings for each method. Shampoo has higher time complexity than PSGD-Affine as the most expensive operations of Shampoo, i.e., eigenvalue decompositions (EVD) for calculating the matrix fourth root inverse in (47), are less efficient than the most expensive operations of the PSGD-Affine, i.e., the backward substitutions in (41), on modern graphic cards. Regarding the test classification error rates, the mean numbers averaged over ten runs are $0.95\%$, $0.8\%$ and $0.7\%$ for SGD, Shampoo and PSGD-Affine, respectively.
Figure 3: (a) Train loss versus epochs for the LeNet5 benchmark. (b) The same as (a) except that the x-axis shows the wall-time running on a RTX 2080 Nvidia graphic card.
IV-B Comparison with L-BFGS
We consider the same task as the above experiment, but replace LeNet5 with a logistic regression model using features $\{1,x_{i,j},x_{i_{1},j_{1}}x_{i_{2},j_{2}}\}$, where $x_{i,j}$, $1\leq i,j\leq 28^{2}$, is the pixel value at location $(i,j)$. This results in a large scale logistic regression problem with over $6\times 10^{6}$ coefficients to fit. L-BFGS is the selected method for this type of problem. Since BFGS is sensitive to gradient noises as shown in Fig. 2, we set batch size to the maximum allowable number on our hardware, $500$. Still, L-BFGS has a good chance, roughly $1/3$, to diverge. Fig. 4 shows the typical convergence curves of SGD, L-BFGS and PSGD-LRA with proper hyper-parameter tunings. From Fig. 4, we see that the line search step from L-BFGS not only makes it has a higher time complexity than PSGD-LRA, but also a higher regression loss plateau. Regarding generalization performance, mean test classification error rates over ten runs of L-BFGS and PSGD-LRA are about $2\%$, and $2.3\%$ for SGD.
Figure 4: (a) Train loss versus epochs for the logistic regression benchmark. (b) The same as (a) except that the x-axis shows the wall-time running on a RTX 2080 Nvidia graphic card.
IV-C Preconditioner Fitting with Stochastic Gradients
For the NLL losses from the exponential family, inverses of the Fisher or empirical Fisher information matrices or their square roots are commonly used as the preconditioners due to their close connections to the Hessians. In this experiment, we check the preconditioner estimation with pairs $(v,g)$ and $(v,h)$. From Table I, we see that for the closed-form solutions, $v$ is a dummy variable and does not shows up in the solutions for $P$ since $E_{v}[v^{T}P^{-1}v]={\rm tr}(P^{-1})$. For preconditioner fitting with PSGD using pair $(v,g)$, $v$ is simply an auxiliary variable so that we can reuse the same math and code for fitting $P$ with pair $(v,h)$ to fit with pair $(v,g)$.
Fig. 5 shows the typical convergence curves for a few methods assuming that the Hessian is decomposable. Closed-form solution Shampoo [7] and its recent improved version, Combining AxeS PReconditioners (CASPR) [9], can only roughly approximate the solutions and perform the worse. The misadjustments of PSGD-Affine with pair $(v,g)$ cannot reduce to zero with a constant finite step size, which is $\mu=0.1$ here. Still, it significantly outperforms Shampoo like coarse approximations. PSGD-Affine with pair $(v,h)$ performs the best, and converges to $H^{-1}$ with errors only bounded by the machine precisions.
Figure 5: Stochastic gradients and Hessian-vector products for preconditioner fitting, where $H\in\mathbb{R}^{100\times 100}$ is decomposable as $H=H^{\prime}\otimes H^{\prime}$ with $H^{\prime}_{i,i}=1$ and $H^{\prime}_{i,j}=0.5$ for $|i-j|=1$.
V Conclusion
In this paper, we have studied a wide range of stochastic Hessian fitting methods using stochastic gradients or Hessian-vector products. Exact closed-form solutions are available only for a few simple cases like a diagonal Hessian and a dense one. They do converge sublinearly to the optimal solution, but may involve numerically challenging matrix operations like inverses and eigenvalue decompositions. The stochastic Hessian fitting criterion from the preconditioned stochastic gradient descent (PSGD) method is thoroughly analyzed in the Euclidean space, the manifold of symmetric positive definite (SPD) matrices and a variety of Lie groups. It is shown to be strongly convex under mild conditions only on a specific Lie group with one possible local coordinate definition, see the Proposition 5 in Section III.A for details. This unique discovery facilitates our designs of several stochastic Hessian fitting methods with linear convergences, self-normalized step sizes, and different time and space complexities, see Table III in Section III.D for a summary. Practically, these Lie group stochastic Hessian fitting methods successfully avoid any numerically problematic operations like large matrix inverse or decompositions. Numerical results have confirmed our theoretical analyses, and demonstrated their efficiency and numerical reliability for real world stochastic optimizations.