### v4

# Stochastic Hessian Fittings with Lie Groups

Abstract
This report studies the fitting of Hessian or its inverse for stochastic optimizations using a Hessian fitting criterion from the preconditioned stochastic gradient descent (PSGD) method, which is intimately related to many commonly used second-order and adaptive gradient optimizers, e.g., BFGS, Gaussian-Newton algorithm, natural gradient descent, AdaGrad, etc. Our analyses reveal the efficiency and reliability differences among a wide range of preconditioner fitting methods, from closed-form to iterative solutions, using Hessian-vector products or stochastic gradients only, with Hessian fittings in the Euclidean space, the manifold of symmetric positive definite (SPL) matrices, to a variety of Lie groups. The most intriguing discovery is that the Hessian fitting itself as an optimization problem is strongly convex under mild conditions in certain general Lie groups. This discovery turns Hessian fitting into a well-behaved Lie group optimization problem and facilitates the designs of highly efficient and elegant Lie group sparse preconditioner fitting methods for large-scale stochastic optimizations.
Index Terms:
Stochastic optimization, stochastic gradient descent (SGD), Hessian-vector product, Lie group, strong convexity.
I Introduction
This report considers the problem of Hessian fitting for second-order stochastic optimizations where the objective functions are typically defined as an expectation, say $f(\theta)=E_{z}[\ell(\theta,z)]$ with $\theta\in\mathbb{R}^{n}$ being the parameters to be optimized, and $z$ a random variable or vector that can be sampled to evaluate an unbiased estimate of $f(\theta)$. We generally assume that $f(\theta)$ is second-order differentiable such that a preconditioner $P\succ 0$, say the inverse of the Hessian of $f(\theta)$ when it is convex, can be used to accelerate the convergence of the following stochastic gradient descent (SGD) iteration,
$\theta_{t+1}=\theta_{t}-\mu_{t}P_{t}\,{\partial f_{t}(\theta)}/{\partial\theta}\left.\right|_{\theta=\theta_{t}}$ (1)
to an optimal solution for $\theta$, where the subscript $t$ is the time index, and $f_{t}(\theta)$ is a sampled noisy estimation of $f(\theta)$ at step $t$. To simplify the notations, let us ignore the time index $t$ when doing so causing no misunderstanding, and use $g={\partial f(\theta)}/{\partial\theta}$ and $h={\partial(v^{T}g)}/{\partial\theta}=Hv$ to denote the stochastic gradient and Hessian-vector product associated with a vector $v\sim\mathcal{N}(0,I)$ at a certain iteration step, respectively, where $H={\partial^{2}f(\theta)}/{(\partial\theta^{T}\partial\theta)}$ is the Hessian matrix and we typically do not assume its definiteness. Then, with the following preconditioner estimation criterion proposed in [17],
$c(P;v,h)=h^{T}Ph+v^{T}P^{-1}v$ (2)
we can derive most of the commonly used preconditioners as listed in Table I [18] by solving for $P$ from $\partial{c}/\partial{P}=0$.
  1. 1.
    When the curvature condition $v^{T}h>0$ is met, solution $Phh^{T}P=vv^{T}$ reduces to the secant equation $Ph=v$ , which is used for deriving the classic BFGS and limited-memory BFGS (L-BFGS) like quasi-Newton methods via iterative Hessian fittings [1]. These off-the-shelf methods are not popular for stochastic optimizations as they do not take the gradient noises into consideration.
  2. 2.
    Closed-form solution $P=(H^{2})^{-0.5}$ reduces to Newton’s method when $H\succ 0$, i.e., $P=H^{-1}$. Diagonal form of this solution is particularly popular in machine learning due to its simplicity, e.g., ESGD [12], AdaHessian [13], and Sophia [14] to name a few.
  3. 3.
    When the target is to minimize a negative logarithm likelihood (NLL) loss and $g_{z}={\partial\ell(\theta,z)}/{\partial\theta}$ is a per-sample gradient for each $z$, closed-form solution $P^{-2}=E_{z}[g_{z}g_{z}^{T}]$ reduces to the (empirical) Fisher information matrix, whose inverse is used in the Gaussian-Newton and natural gradient methods [2, 3] as preconditioners. Its Kronecker factorized approximated solution is popular for the learning of matrix parameters [5, 6, 2].
  4. 4.
    Closed-form solution $P_{t}=\left(\sum_{i=1}^{t}g_{i}g_{i}^{T}\right)^{-0.5}$ reduces to the preconditioner used in the AdaGrad [4], a large family of optimizers including many popular examples, e.g., the RMSProp [10] and Adam(W) [11] that use diagonal preconditioners, and a few methods with the name Shampoo the most well known [6, 7, 8, 9] that use Kronecker factorized approximate closed-form solutions.
Except for the trivial case of diagonal preconditioners, all the last three closed-form solutions for $P$ in Table I require matrix or matrix square root inverse operations, which could cause numerical difficulties with finite precision arithmetic [8]. Furthermore, exact closed-form solutions are seldom available for sparse Hessian estimations. The iterative Hessian fitting method of BFGS requires that $v^{T}h$ is well above zero, which can be hardly met in stochastic optimizations. The preconditioned stochastic gradient descent (PSGD) method fits $P$ iteratively in a variety of Lie groups [17, 18, 19], and have successfully avoided the numerical difficulties of those closed-form solutions and also lift the limitations of BFGS. Recent empirical results have further confirmed PSGD’s excellent performance on a handful of large scale machine learning related stochastic optimizations [20, 21]. Still, there lacks a deeper theoretical understanding of the pros and cons of the preconditioners listed in Table I, disregarding their certain practical limitations like computational complexities and numerical difficulties with finite precision arithmetic. One major contribution of this report is to show that the problem of Hessian fitting by minimizing criterion (2) converges linearly to the optimal solution in the manifold of symmetric positive definite (SPD) matrices, and furthermore, is strongly convex in certain Lie groups, see Proposition 5 in Section III.A, while those closed-form solutions only converge sublinearly. Another significant contribution is the invention of a L-BFGS like low-rank approximation (LRA) version of PSGD  [^1] This LRA preconditioner fitting work has been presented in the NeurIPS HOOML workshop, 2022 [19].. Lastly, the insights brought up by the studies here also help to improve previous PSGD implementations [17, 18, 19] in several aspects.
The notations in this paper are fairly standard. For example, $A\succ 0$ means that $A$ is a SPD matrix, $E_{v}[\cdot]$ takes expectation with respect to a certain random variable $v$, $\mathcal{N}(0,I)$ denotes the standard normal distribution, $A^{T}$ is the transpose of $A$, $A^{0.5}$ denotes the principal square root of $A$, $A^{-2T}$ is the shorthand for $(A^{-2})^{T}$, $\|A\|_{2}$ is the spectral norm of $A$, ${\rm tr}(A)$ takes the trace of $A$, $\lambda_{\min}(A)$ takes the minimum eigenvalue of $A$, ${\rm vec}(A)$ vectorizes matrix $A$ by stacking its columns successively, GL$(n,\mathbb{R})$ stands for the general linear group of $n\times n$ invertible matrices, and $\otimes$ denotes Kronecker product. We say that $A$ and $B$ commute if $[A,B]=AB-BA=0$. For a symmetric matrix $A\in\mathbb{R}^{n\times n}$, we only need to vectorize either its upper or lower triangular part, and denote this operation as ${\rm vech}(A)$. Clearly, ${\rm vec}(A)=S\,{\rm vech}(A)$, where $S\in\mathbb{R}^{n^{2}\times[0.5n(n+1)]}$ is a sparse full column rank matrix with elements being 00 or $1$. By convention, we use capital letters for matrices, and lower case letters for vectors and scalars.
TABLE I: Variations of criterion (2) for preconditioner fitting, where $v$ is a dummy variable in 2), 3) and 4) as $E_{v}[v^{T}P^{-1}v]={\rm tr}(P^{-1})$.
Criterion Solution for $P$
1) $c(P;v,h)$ $Phh^{T}P=vv^{T}$
2) $E_{v}[c(P;v,h)]$ $P^{-2}=H^{2}$
3) $E_{v,z}[c(P;v,g_{z})]$ $P^{-2}=E_{z}[g_{z}g_{z}^{T}]$
4) $\sum_{i=1}^{t}E_{v_{i}}\left[c(P;v_{i},g_{i})\right]$ $P_{t}^{-2}=\sum_{i=1}^{t}g_{i}g_{i}^{T}$
II Candidates for Stochastic Hessian Fitting
II-A The Problem Setups
We are to fit the Hessian or its inverse sequentially, i.e., updating the Hessian estimate once given every pair of vector and its associated Hessian-vector product, $(v,h)$, with model,
$h=Hv+\epsilon$ (3)
where $H\in\mathbb{R}^{n\times n}$ is the unknown Hessian matrix to be estimated, and $\epsilon$ denotes any possible modeling errors, e.g., truncation and quantization errors when the Hessian-vector product $h$ is approximated with a finite-difference method, or the randomness error when a stochastic yet unbiased proxy of the target function $f(\theta)=E_{z}[\ell(\theta,z)]$ is adopted for Hessian-vector product evaluation, or the structural errors caused by using a certain form of sparse matrices to fit dense Hessians. We assume that $\epsilon$ is a random error term independent of $v$.
We choose to update $P$ by minimizing the preconditioner fitting criterion (2), whose expectation with respect to $v$ is,
$c(P)=$ $E_{v\sim\mathcal{N}(0,I)}[c(P;v,h)]$
$=$ ${\rm tr}\left(PE[hh^{T}]+P^{-1}E[vv^{T}]\right)$
$=$ ${\rm tr}\left(P\left(H^{2}+E[\epsilon\epsilon^{T}]\right)+P^{-1}\right)$ (4)
Then, the optimal SPD solution minimizing $c(P)$ can be shown to be [17]
$P=\left(H^{2}+E[\epsilon\epsilon^{T}]\right)^{-0.5}$ (5)
by solving equation ${\partial c(P)}/{\partial P}=0$ for $P$. One also can replace the $P$ in (2) with $P^{-1}$ to obtain another criterion for fitting the Hessian directly. Here, we only consider the fitting criterion (2) since the intention is to use $P$ as a preconditioner for accelerating the SGD iteration in (1).
As revealed by (5), the error term $\epsilon$ in (3) helps to damp the preconditioner estimation and has been shown to regularize the optimizer in a perfectly balanced way [17]. Direct use of $(H^{2})^{-0.5}$ as a preconditioner could severely amplify the gradient noises, leading to divergence [17]. Definiteness of the Hessian $H$ is important for the optimization of $f(\theta)$, but not relevant to our setups since both (4) and (5) suggest that the signs of $\lambda(H)$ do not matter to the finial solutions for $P$. Hence, without loss of generality, we always assume that $H$ is a constant positive definite matrix in our performance studies. Just bear in mind that except for the BFGS, other preconditioners derived from (2) are always valid regardless of the definiteness of $H$.
We will further ignore the term $\epsilon$ most of the time to simplify our performance analyses. Still, note that $c(P)$ always has form $c(P)={\rm tr}(PA+P^{-1})$, where $A=H^{2}+E[\epsilon\epsilon^{T}]$ with pair $(v,h)$, and $A=E_{z}[g_{z}g_{z}^{T}]$ or $\sum_{i=1}^{t}[g_{i}g_{i}^{T}]$ with pair $(v,g)$. Thus, properties proved for one case can be easily transferred to the other cases.
Criterion (2) is a well defined function for any $P$ with $\det(P)\neq 0$, i.e., the group GL$(n,\mathbb{R})$. For our purpose, we further require $P\succ 0$ such that it is a valid preconditioner for the SGD iteration in (1). The implicitly implied symmetry constraint, i.e., $P=P^{T}$, will slightly complicate the calculations of the derivatives of $c(P;v,h)$ with respect to $P$. We will bring up this subtleness explicitly when it matters. From now on, by default, the gradient and Hessian always refer to the ones of $c(P;v,h)$ or $c(P)=E_{v}[c(P;v,h)]$ with respect to $P$. Gradient descent (GD) or SGD also only refers to the iterations for updating $P$, not $\theta$, if not further clarified.
II-B Second Order Approximation of Criterion (2)
Let $dP$ denote a small perturbation around $P$. We are to expand the fitting loss in (2) up to its second order terms of $dP$. To begin with, we expand $(P+dP)^{-1}$ around $P^{-1}$ as,
$\quad dP^{-1}=(P+dP)^{-1}-P^{-1}$
$=(I-P^{-1}dP+P^{-1}dP\,P^{-1}dP+\mathcal{O}((dP)^{3}))P^{-1}-P^{-1}$
$=-P^{-1}dP\,P^{-1}+P^{-1}dP\,P^{-1}dP\,P^{-1}+\mathcal{O}((dP)^{3})$ (6)
where $\mathcal{O}((dP)^{3})$ includes the third and higher order terms of $dP$ that are negligible when $\|dP\|$ is small enough. Now, with (6), we can expand (2) as below,
$dc(P;v,h)=$ $h^{T}dP\,h+v^{T}dP^{-1}\,v$
$=$ $h^{T}dP\,h-v^{T}P^{-1}dP\,P^{-1}v$
$+v^{T}P^{-1}dP\,P^{-1}dP\,P^{-1}v$ (7)
where we have suppressed the term $\mathcal{O}((dP)^{3})$ to simplify our equation. Clearly, the $dc(P;v,h)$ in (7) becomes a quadratic function of $dP$ after truncating its higher order terms. To maximize the reduction of $c(P;v,h)$, we can solve for the optimal $dP$ by letting $dc(P;v,h)/dP=0$, i.e., the following equation,
$hh^{T}-P^{-T}vv^{T}P^{-T}+P^{-T}vv^{T}P^{-T}dP^{T}\,P^{-T}$
$+P^{-T}dP^{T}\,P^{-T}vv^{T}P^{-T}=0$ (8)
where we do not assume $P=P^{T}$ yet to simplify the calculation of derivatives. From (8), it is ready to identify the gradient of $c(P;v,h)$ with respect to $P$ as below,
$\frac{\partial c(P;v,h)}{\partial P}=hh^{T}-{P}^{-T}vv^{T}{P}^{-T}$ (9)
Note that (9) gives the steepest ascent direction of $c(P;v,h)$ in the space $\mathbb{R}^{n\times n}$. With constraint $P=P^{T}$, we should double the off-diagonal elements of the gradient in (9) to obtain the steepest ascent direction in the space $\mathbb{R}^{0.5n(n+1)}$. We can choose either form for the update of $P$ with GD or SGD. But, we do need to calculate the Hessian with respect to $P$ in the space $\mathbb{R}^{0.5n(n+1)}$. We typically take the expectation of criterion (2), i.e., the $c(P)=E_{v\sim\mathcal{N}(0,I)}[c(P;v,h)]$ given in (4), for the performance analyses. There is the following convex property of $c(P)$.
Proposition 1: Criterion $c(P)$ is a convex function of $P$ in set $\{P|P\succ 0\}$.
Proof: From (7), we can rewrite the second order term of $dc(P)$ as
$\quad E_{v\sim\mathcal{N}(0,I)}\left[v^{T}P^{-1}dP\,P^{-1}dP\,P^{-1}v\right]$
$={\rm tr}\left(E_{v}\left[P^{-1}dP\,P^{-1}dP\,P^{-1}vv^{T}\right]\right)$
$={\rm tr}\left(P^{-1}dP\,P^{-1}dP\,P^{-1}\right)$
$=\left({\rm vec}\left(\left(P^{-1}dP\right)^{T}\right)\right)^{T}{\rm vec}\left(P^{-1}dP\,P^{-1}\right)$
$=\left({\rm vec}\left(dP\,P^{-1}\right)\right)^{T}{\rm vec}\left(P^{-1}dP\,P^{-1}\right)$
$=\left(\left(P^{-1}\otimes I\right){\rm vec}(dP)\right)^{T}\left(\left(P^{-1}\otimes P^{-1}\right){\rm vec}(dP)\right)$
$=\left({\rm vec}(dP)\right)^{T}\left(P^{-1}\otimes I\right)\left(P^{-1}\otimes P^{-1}\right){\rm vec}(dP)$
$=\left({\rm vec}(dP)\right)^{T}\left(P^{-2}\otimes P^{-1}\right){\rm vec}(dP)$
$=\left({\rm vech}(dP)\right)^{T}S^{T}\left(P^{-2}\otimes P^{-1}\right)S\,{\rm vech}(dP)$ (10)
where we have used the relation ${\rm tr}(A^{T}B)$$=$$\left({\rm vec}(A)\right)^{T}{\rm vec}(B)$ to arrive the fourth line from the third one, ${\rm vec}(ABC)=(C^{T}\otimes A){\rm vec}(B)$ from the fifth to the sixth line, $(AB)\otimes(CD)=(A\otimes C)(B\otimes D)$ from the third to the last to the next line, and definition ${\rm vec}(A)=S\,{\rm vech}(A)$ for $A\succ 0$ to get the last line. Since $P^{-2}\otimes P^{-1}\succ 0$ and $S$ has full column rank, the Hessian with respect to $P$, i.e.,
$S^{T}\left(P^{-2}\otimes P^{-1}\right)S$
due to the last line of (10), is positive definite. $\square$
Note that following the proof of Proposition 1, we also can show that $c(P;v,h)$ is convex as well. Actually, this convexity property is independent of the definiteness of $H$. In general, the quadratic term in (7) can be negative without the implicit constraint $dP=dP^{T}$ even when $P\succ 0$. Thus, the symmetry constraint is necessary for the convexity of $c(P)$.
II-C Hessian Fitting in the Euclidean Space
II-C1 Gradient Descent
One obvious choice is to update $P$ with the gradient given in (9). Let us start from an initial guess $P_{0}\succ 0$, and update $P$ at step $t$ with SGD as
$P_{t+1}=P_{t}-\mu(h_{t}h_{t}^{T}-{P}_{t}^{-1}v_{t}v_{t}^{T}{P}_{t}^{-1})$ (11)
where $\mu$ is a small enough positive step size, $(v_{t},h_{t})=(v_{t},Hv_{t})$ with $v_{t}\sim\mathcal{N}(0,I)$ is a pair of vector and its associated Hessian-vector product at step $t$. Note that $P_{t}$ is always symmetric as long as $P_{0}$ is.
From the proof of Proposition 1, we see that the Hessian of $c(P)$ with respect to $P$ is neither lower nor upper tightly bounded. Thus, it is generally not easy to characterize the convergence rate of the iteration in (11). Still, following the regret analysis [4, 11], we can show that with a proper constant step size $\mu$, the sequence given by (11) converges sublinearly to $H^{-1}$ when $H\succ 0$. We only sketch the key steps here. Assuming that $\mu$ is small enough such that we always have $P_{t}\succ 0$ as long as $P_{0}\succ 0$. Then, via the convexity of $c(P;v,h)$, we can show that
$\quad c(P_{t};v_{t},h_{t})-c(H^{-1};v_{t},h_{t})$
$\leq\frac{\left\|P_{t}-H^{-1}\right\|_{F}^{2}-\left\|P_{t+1}-H^{-1}\right\|_{F}^{2}}{2\mu}+\frac{\mu}{2}\left\|\frac{\partial c(P_{t};v_{t},h_{t})}{\partial P}\right\|_{F}^{2}$
Then, we take expectation on both sides of the above equation with respect to $v$, and sum over the time index $t$ to have
$c\left(\sum_{i=1}^{t}P_{i}/t\right)-c(H^{-1})=\mathcal{O}(1/\sqrt{t})$
by setting $\mu\propto 1/\left(\sqrt{t}\,\max_{1\leq i\leq t}E_{v_{i}}\left[\left\|\frac{\partial c(P_{i};v_{i},h_{i})}{\partial P}\right\|_{F}\right]\right)$. Still, the regret analysis does not exclude possible better convergence results of (11) with time-varying step sizes.
II-C2 Newton’s Method
Another choice is to take the expectation on the left side of (8) to have
$H^{2}-P^{-2}+P^{-2}dP\,P^{-1}+P^{-1}dP\,P^{-2}=0$ (12)
and then solve for the optimal $dP$ either directly as
$\>\>{\rm vec}(dP)=\left(P^{-1}\otimes P^{-2}+P^{-2}\otimes P^{-1}\right)^{-1}{\rm vec}\left(P^{-2}-H^{2}\right)$
$=\left(I\otimes P^{-1}+P^{-1}\otimes I\right)^{-1}(P\otimes P){\rm vec}\left(P^{-2}-H^{2}\right)$
$=\left(I\otimes P^{-1}+P^{-1}\otimes I\right)^{-1}{\rm vec}\left(I-PH^{2}P\right)$ (13)
or via the matrix sign function iteratively [22], which is clear after rewriting (12) into the following familiar continuous Lyapunov function form,
$P^{-1}dP+dP\,P^{-1}=I-PH^{2}P$
where $H^{2}$ is given or estimated from $H^{2}=E[hh^{T}]$ by replacing $E[\cdot]$ with sample averages. Eq. (13) or the method in [22] gives a numerically stable but costly solution for $dP$.
But, if we assume that $P$ and $dP$ commute such that $P^{-2}dP\,P^{-1}=P^{-1}dP\,P^{-2}$, then (12) immediately gives the optimal $dP$ as below,
$dP=-0.5P(H^{2}-P^{-2})P^{2}=0.5(P-PH^{2}P^{2})$ (14)
Now, (14) suggests the following Newton’s method for updating $P$,
$P_{t+1}=1.5P_{t}-0.5P_{t}H^{2}P_{t}^{2}$ (15)
Proposition 2: Starting from an initial guess $P_{0}\succ 0$ that commutes with $H$ and satisfies condition $20<\lambda(HP)<(\sqrt{17}-1)/$ 2, the Newton’s method in (15) converges to $H^{-1}$ quadratically.
Proof: First, let us show that $P_{t+1}$ commutes with $H$ if $P_{t}$ does. With (15), we have
$P_{t+1}H-HP_{t+1}$
$=$ $1.5(P_{t}H-HP_{t})-0.5(P_{t}H^{2}P_{t}^{2}H-HP_{t}H^{2}P_{t}^{2})$
$=$ $-0.5(P_{t}H^{3}P_{t}^{2}-P_{t}H^{3}P_{t}^{2})=0$
Hence, $P_{t}$ always commutes with $H$ as long as $P_{0}$ does. Also, we can easily show that $P_{t}$ is symmetric as long as $P_{0}$ is.
Now, let us define $R_{t}=HP_{t}-I$. Similarly, we can show that $P_{t}$, $H$ and $R_{t}$ all commute with each other and are symmetric. Then, by (15), we have
$R_{t+1}=$ $1.5HP_{t}-0.5HP_{t}H^{2}P_{t}^{2}-I$
$=$ $1.5HP_{t}-0.5(HP_{t})^{3}-I$
$=$ $1.5(R_{t}+I)-0.5(R_{t}+I)^{3}-I$
$=$ $-0.5(R_{t}+3I)R_{t}^{2}$ (16)
We require
$\|0.5(R_{t}+3I)R_{t}\|_{2}<1$
to make (16) a contraction mapping such that $\|R_{t+1}\|_{2}<\|R_{t}\|_{2}$, which suggests that the eigenvalues of $P_{t}H$ should be bounded by
$0<\lambda(P_{t}H)<\frac{\sqrt{17}-1}{2}$ (17)
Then, with the bounds given in (17) and the last line of (16), we can show the following quadratic convergence,
$\frac{\|R_{t+1}\|_{2}}{\|R_{t}\|_{2}^{2}}=0.5\|R_{t}+3I\|_{2}<\frac{\sqrt{17}+3}{4}$
Note that the lower bound of $\lambda(P_{t}H)$ in (17) already implies that $H$ cannot be singular such that $H^{-1}$ exists. Also please note that for the rate of quadratic convergence, it is sufficient to show that ${\|R_{t+1}\|_{2}}/{\|R_{t}\|_{2}^{2}}$ is tightly upper bounded, not necessarily strictly upper bounded by $1$square
It is trivial to choose a starting point $P_{0}$ that commutes with $H$, e.g., a $P_{0}\propto I$. Indeed, starting from a $P_{0}\propto I$, we see that $P_{t}$ is a linear combination of the terms in $\{I,H^{2},H^{8},\ldots\}$ and thus commutes with $H$. However, in practice, the Newton’s method given in (15) is vulnerable to round-off errors that eventually break the commuting property between $P_{t}$ and $H$.
Another related method is the Newton-Schulz iteration [23], which is numerically stable, and more often used to calculate $(H^{2})^{-0.5}$ iteratively. We seldom use these Newton’s methods in practice. We present them here mainly for theoretical interests, especially their connections to the iterative methods for matrix square root calculations [22, 23].
II-C3 A Few Closed-Form Solutions
For the sake of completeness, let us also briefly discuss the convergence rate of a few closed-form solutions in the Euclidean space. The most straightforward and commonly used choice is $P_{t}$$=$$\left(\left(P_{0}^{-2}+\sum_{i=1}^{t}h_{i}h_{i}^{T}\right)/(t+1)\right)^{-0.5}$, e.g., methods from the AdaGrad family [4], which can be put in a fancier form as below
$P_{t+1}=\left(\frac{t+1}{t+2}P_{t}^{-2}+\frac{1}{t+2}h_{t}h_{t}^{T}\right)^{-0.5}$ (18)
Since $h_{t}=Hv_{t}$ and $v_{t}\sim\mathcal{N}(0,I)$, we have
$P_{t}=$ $\left(H\left(P_{0}^{-2}/(t+1)+\sum_{i=1}^{t}v_{i}v_{i}^{T}/(t+1)\right)H\right)^{-0.5}$
$=$ $\left(H\left(I+\mathcal{O}(1/t)\right)H\right)^{-0.5}$
$=$ $H^{-1}+\mathcal{O}(1/t)$
Hence, the closed-form solution given by (18) converges to $H^{-1}$ sublinearly with rate $\mathcal{O}(1/t)$ when $H\succ 0$.
Another closed-formed solution can be obtained from (8) by solving for $P$ directly after setting $dP=0$. Specifically, (8) suggests that a stationary point of $P$ satisfies
$P\left(\sum_{i=1}^{t}h_{i}h_{i}^{T}\right)P-\sum_{i=1}^{t}v_{i}v_{i}^{T}=0$ (19)
Then, we can directly solve the above Riccati equation for $P$ when both $\sum_{i=1}^{t}v_{i}v_{i}^{T}$ and $\sum_{i=1}^{t}h_{i}h_{i}^{T}$ are not singular, see the proof of Proposition 3 in [17]. Clearly, (19) reduces to (18) for $t\rightarrow\infty$. But, this solution is exact even with a finite $t$ as long as the solution is unique. Similar to Newton’s methods, this closed-form solution is mainly for theoretical interests.
One more family of solutions worth mentioning here are the BFGS like methods that modify $P_{t}$ with certain rank-2 updates such that $P_{t+1}h_{t}=v_{t}$ [1]. They are essentially iterative methods for Hessian fittings, but with a few heuristic closed-form update rules for $P$. These methods have two major limitations: the pair $(v_{t},h_{t})$ needs to satisfy the curvature condition $v_{t}^{T}h_{t}>0$; they diverge easily with noisy Hessian-vector products as inputs. Note that generally, these heuristic methods could diverge, or might not converge to $\left(H^{2}\right)^{-0.5}$.
II-D Hessian Fitting in the Manifold of SPD
One possible way to define a local coordinate for $P$ in the manifold of SPD is to let $dP$ be
$dP=P\mathcal{E}+\mathcal{E}P^{T}$ (20)
where $\mathcal{E}$ is a sufficiently small matrix, but not necessarily symmetric in general. Clearly, $dP$ is symmetric if $\mathcal{E}$ is, regardless of the symmetry property of $P$. On the other hand, we can show that $\mathcal{E}$ is symmetric if $P\succ 0$.
Proposition 3: Eq. (20) induces a new local coordinate system for the manifold $\{P|P\succ 0\}$.
Proof: We are to show that (20) defines a bijective mapping between $dP$ and $\mathcal{E}$ when $P\succ 0$, which is partially done after we vectorize both sides of (20) to connect $dP$ to $\mathcal{E}$ as,
${\rm vec}(\mathcal{E})=(I\otimes P+P\otimes I)^{-1}{\rm vec}(dP)$ (21)
Still, we need to show that (21) preserves the symmetry properties of $dP$ and $\mathcal{E}$. Clearly, a symmetric $\mathcal{E}$ maps to a symmetric $dP$ by (20). The left part is to show that (21) also maps a symmetric $dP$ to a symmetric $\mathcal{E}$. To begin with, let us introduce a permutation matrix $\Pi$ such that ${\rm vec}(A^{T})=\Pi\,{\rm vec}(A)$. Then, for any matrix $X$ with the same size as $P$, we have
$(P\otimes I+I\otimes P)\Pi\,{\rm vec}(X)=$ $(P\otimes I+I\otimes P){\rm vec}(X^{T})$
$=$ ${\rm vec}(X^{T}P+PX^{T})$
$=$ ${\rm vec}\left((PX+XP)^{T}\right)$
$=$ $\Pi{\rm vec}(PX+XP)$
$=$ $\Pi(I\otimes P+P\otimes I){\rm vec}(X)$
Since $X$ is arbitrary, the above equation suggests that
$\Pi^{-1}(P\otimes I+I\otimes P)=(I\otimes P+P\otimes I)\Pi^{-1}$ (22)
Then, from (21), we have
${\rm vec}(\mathcal{E}^{T})=$ $\Pi\,{\rm vec}(\mathcal{E})$
$=$ $\Pi(I\otimes P+P\otimes I)^{-1}{\rm vec}(dP)$
$=$ $\left((I\otimes P+P\otimes I)\Pi^{-1}\right)^{-1}{\rm vec}(dP)$
$=$ $\left(\Pi^{-1}(I\otimes P+P\otimes I)\right)^{-1}{\rm vec}(dP)$
$=$ $(I\otimes P+P\otimes I)^{-1}\Pi{\rm vec}(dP)$
$=$ $(I\otimes P+P\otimes I)^{-1}{\rm vec}(dP^{T})$
$=$ $(I\otimes P+P\otimes I)^{-1}{\rm vec}(dP)$
$=$ ${\rm vec}(\mathcal{E})$
where we have used (22) from the third to the fourth line. Thus, we have $\mathcal{E}=\mathcal{E}^{T}$ when $dP=dP^{T}$. Finally, note that the condition $P\succ 0$ not only implies $dP=dP^{T}$, but also makes (21) bijective. This completes the proof. $\square$
By Proposition 3, (20) induces a new coordinate system for the manifold of SPD matrices. For a positive scalar $P$, integration on the right side of (21) gives the closed-form solution for transforming $P$ to the new one, i.e., $0.5\log P+{\rm constant}$. In general, we have no need to write down this transform explicitly for a matrix $P$, and (20) is enough for our purpose.
Now, we can rewrite the $dc(P;v,h)$ in (7) in the new coordinate as below after ignoring term $\mathcal{O}(\mathcal{E}^{3})$,
$\quad dc(P;v,h)$
$=h^{T}(P\mathcal{E}+\mathcal{E}P^{T})h-v^{T}P^{-1}(P\mathcal{E}+\mathcal{E}P^{T})P^{-1}v$
$\quad+v^{T}P^{-1}(P\mathcal{E}+\mathcal{E}P^{T})P^{-1}(P\mathcal{E}+\mathcal{E}P^{T})P^{-1}v$
$=h^{T}P\mathcal{E}h+h^{T}\mathcal{E}Ph-v^{T}\mathcal{E}P^{-1}v-v^{T}P^{-1}\mathcal{E}v+v^{T}\mathcal{E}^{2}P^{-1}v$
$\quad+v^{T}\mathcal{E}P^{-1}\mathcal{E}v+v^{T}P^{-1}\mathcal{E}P\mathcal{E}P^{-1}v+v^{T}P^{-1}\mathcal{E}^{2}v$ (23)
where in the last line we have assumed that $P=P^{T}$. Again, (23) is a quadratic function of $\mathcal{E}$, and we can solve for the $\mathcal{E}$ that reduces $dc(P;v,h)$ the most from the following equation,
$\quad Phh^{T}+hh^{T}P-vv^{T}P^{-1}-P^{-1}vv^{T}+vv^{T}P^{-1}\mathcal{E}^{T}$
$+\mathcal{E}^{T}vv^{T}P^{-1}+vv^{T}\mathcal{E}^{T}P^{-1}+P^{-1}\mathcal{E}^{T}vv^{T}$
$+P^{-1}vv^{T}P^{-1}\mathcal{E}^{T}P+P\mathcal{E}^{T}P^{-1}vv^{T}P^{-1}$
$+P^{-1}vv^{T}\mathcal{E}^{T}+\mathcal{E}^{T}P^{-1}vv^{T}=0$ (24)
Then, we can identify the gradient in the new coordinate from (24), and update $P$ with SGD given pair $(v_{t},h_{t})$ at step $t$ as,
$\mathcal{E}_{t}=$ $-\mu(P_{t}h_{t}h_{t}^{T}+h_{t}h_{t}^{T}P_{t}-v_{t}v_{t}^{T}P_{t}^{-1}-P_{t}^{-1}v_{t}v_{t}^{T})$
$P_{t+1}=$ $P_{t}+(P_{t}\mathcal{E}_{t}+\mathcal{E}_{t}P_{t})$ (25)
Clearly, with (25), $P_{t}$ always is symmetric as long as $P_{0}$ is.
Proposition 4: Starting from an initial guess $P_{0}\succ 0$ that commutes with $H$ and with a constant step size $0<\mu<0.5/\max_{t}{\lambda_{\max}(H+H^{2}P_{t})}$, the sequence $P_{t}$ given by (25) converges to $H^{-1}$ linearly with rate $1-8\mu\,\lambda_{\min}(H)$.
Proof: We take expectation on both sides of (25) to have
$\quad E[P_{t+1}]$
$=E[P_{t}]-\mu\left(E[P_{t}^{2}]H^{2}+2E[P_{t}H^{2}P_{t}]+H^{2}E[P_{t}^{2}]-4I\right)$
$=E[P_{t}]-\mu\left(\left(E[P_{t}]\right)^{2}H^{2}+2E[P_{t}]H^{2}E[P_{t}]\right.$
$\qquad\left.+H^{2}\left(E[P_{t}]\right)^{2}-4I+\mathcal{O}(\left(P_{t}-E[P_{t}]\right)^{2})\right)$ (26)
where $E[\cdot]$ is a shorthand for $E_{v_{1}\sim\mathcal{N}(0,I),\ldots,v_{t}\sim\mathcal{N}(0,I)}[\cdot]$. The term $\mathcal{O}(\left(P_{t}-E[P_{t}]\right)^{2})$ in the last line of (26) is negligible with a small $\mu$. We further drop the notation $E[\cdot]$ in (26), and just use $P_{t}$ to denote its expectation to simplify our writings.
First, we are to show that $P_{t+1}$ commutes with $H$ if $P_{t}$ does. Assuming $P_{t}H=HP_{t}$, we can simplify (26) to
$P_{t+1}=P_{t}-4\mu(H^{2}P_{t}^{2}-I)$ (27)
Then, it is not difficult to show that $P_{t+1}$ also commutes with $H$, similar to the proofs in Proposition 2. Thus, $P_{t}$ always commutes with $H$ as long as $P_{0}$ does.
Second, let us define fitting error $R_{t}=HP_{t}-I$ and show that $\|R_{t}\|_{2}\rightarrow 0$ for $t\rightarrow\infty$. Clearly, $H$, $P_{t}$ and $R_{t}$ all are symmetric and commute with each other. From (27), we have
$R_{t+1}=$ $HP_{t+1}-I$
$=$ $HP_{t}-4\mu H\left((HP_{t})^{2}-I\right)-I$
$=$ $(R_{t}+I)-4\mu H\left((R_{t}+I)^{2}-I\right)-I$
$=$ $(I-8\mu H-4\mu HR_{t})R_{t}$
$=$ $(I-4\mu H-4\mu H^{2}P_{t})R_{t}$ (28)
From the last line of (28), we see that $\mu$ needs to satisfy
$0<\mu<\frac{1}{2\lambda_{\max}(H+H^{2}P_{t})}$
to ensure the monotonic convergence of $P_{t}$. Noting that $\|R_{t}\|_{2}\rightarrow 0$ for $t\rightarrow\infty$ and $[H,R_{t}]=0$, the second to the last line of (28) gives the following asymptotic convergence rate,
$1>\frac{\|R_{t+1}\|_{2}}{\|R_{t}\|_{2}}=\|I-8\mu H\|_{2}\geq 1-8\mu\,\lambda_{\rm min}(H)$
Thus, $P_{t}$ converges linearly to $H^{-1}$ when $\lambda_{\rm min}(H)>0$. $\square$
Again, similar to the proof of Proposition 2, the condition that $P$ and $H$ commute simplifies the proof, but could be too strict to be satisfied in practice. Yet, the linear convergence rate $1-8\mu\,\lambda_{\min}(H)$ matches well with observations.
Compared with SGD in the Euclidean space, linear convergence of the iterations given by (25) with a small enough constant step size is a good sign of progress. Intuitively, from (20), we see that $dP$ is roughly proportional to $P$. Hence, (20) mimics a Euclidean space varying step size adaptation. Following a similar way for the proof of Proposition 1, from (23), we can show that the Hessian of $c(P)$ in the new coordinates is
$S^{T}\left[3(P^{-1}\otimes I)+P^{-2}\otimes P\right]S$
which again is neither lower nor upper tightly bounded. Thus, we have moved the needle a little bit, but not that much.
It is interesting to check the Newton’s methods in the new coordinates as well. Following the developments of the Newton’s methods in the Euclidean space, we can take expectation on the left side of (24) and solve for the optimal $\mathcal{E}$ directly. Generally, the two sets of Newton’s methods are different as the coordinate transform defined by (20) are not linear. However, if we assume that $\mathcal{E}$ and $P$ commute, we can readily show that $dP$ and $P$ commute as well, and eventually the two Newton’s methods are the same in this case.
II-E Hessian Fitting with Lie Groups
Linear convergence for Hessian fitting in Lie groups will not be a surprise since a Lie group is just a special manifold equipped with extra algebra properties that could further facilitate the development of Hessian fitting methods. One significant advantage of fitting $P$ with certain Lie groups is that they enable us to tightly lower bound the Hessian of $c(P)$. We leave these developments to the next section.
II-F Summary of the Hessian Fitting Methods
We summarize most of the above candidates for Hessian fitting in Table II. To confirm our analyses, Fig. 1 shows a set of typical convergence curves when $H$ is a small Hilbert matrix. Except for Newton method that starts from $P_{0}=0.02I$, all the other methods start from $P_{0}=I$. The step sizes for SGD in the Euclidean space and the manifold of SPD matrices are tuned to achieve the maximum rate of convergence. Step sizes for SGD with Lie group and Newton’s method are already normalized to $1$. Except for the SGD (11) in the Euclidean space and the closed-form solution (18) that converge sublinearly, all the other methods converge to $H^{-1}$ within reasonable numbers of iterations. These empirical convergence rates match well with the predictions in Table II.
TABLE II: Candidates for Hessian fitting. Note that $10<\rho<$ 1 is not necessarily the same for different methods.
Method Equation Space Error at $t$
SGD (11) Euclidean $\mathcal{O}(1/\sqrt{t})$
Closed-form (18) Euclidean $\mathcal{O}(1/t)$
SGD (25) SPD $\mathcal{O}(\rho^{t})$
SGD (37) Lie group $\mathcal{O}(\rho^{t})$
Newton (15) Euclidean/SPD $\mathcal{O}(\rho^{2^{t}})$
Figure 1: (a): Typical convergence curves for the five Hessian fitting methods in Table II when $H$ is a $3\times 3$ Hilbert matrix, i.e., $H_{i,j}=1/(i+j-1)$ with $1\leq i,j\leq n$. (b) and (c): Zooming in of the converge curves of SGD in group GL$(n,\mathbb{R})$ and Newton’s method, respectively, for better visualization.
III Hessian Fitting with Lie Groups
III-A Strong Convexity in Certain Lie Groups
It is still not handy to fit Hessian in the manifold of SPD matrices as we need to constantly guard the positive definiteness of $P_{t}$. Also, per Proposition 4, knowledge of $H$ is required to determine the fitting step sizes. Here, we let
$P=Q^{T}Q$ (29)
[…]
With the Lie group constraint, we let $I+\mathcal{E}=e^{\varepsilon G}$, where $\varepsilon\rightarrow 0$ is a small scalar such that $\varepsilon\|G\|\ll 1$, $G$ is the group’s infinitesimal generator, and $\mathcal{E}=\varepsilon G+\varepsilon^{2}G^{2}/2!+\ldots$ is a small matrix. Then, we can have the following two choices for $dQ$,
${\rm choice\,I:}\quad$ $dQ=e^{\varepsilon G}Q-Q=\mathcal{E}Q$
${\rm choice\,II:}\quad$ $dQ=Qe^{\varepsilon G}-Q=Q\mathcal{E}$ (30)
Correspondingly, the $dP$ in (29) will have the following two possible forms,
${\rm choice\,I:}\quad dP=$ $(Q+\mathcal{E}Q)^{T}(Q+\mathcal{E}Q)-P$
$=$ $Q^{T}(\mathcal{E}+\mathcal{E}^{T}+\mathcal{E}^{T}\mathcal{E})Q$
${\rm choice\,II:}\quad dP=$ $(Q+Q\mathcal{E})^{T}(Q+Q\mathcal{E})-P$
$=$ $\mathcal{E}^{T}P+P\mathcal{E}+\mathcal{E}^{T}P\mathcal{E}$ (31)
The second choice of $dP$ in (31) looks almost the same as that in (20) after ignoring the second order term of $\mathcal{E}$. But, now $\mathcal{E}$ is a small matrix associated with the group generator, and no longer required to be symmetric. Substituting (31) back into (7) and noting that $P^{-1}=Q^{-1}Q^{-T}$, we can arrive at the following two possible choices for $dc(P;v,h)$,
$dc_{I}(P;v,h)=h^{T}Q^{T}(\mathcal{E}+\mathcal{E}^{T})Qh-v^{T}Q^{-1}(\mathcal{E}+\mathcal{E}^{T})Q^{-T}v$
$\qquad\quad+h^{T}Q^{T}\mathcal{E}^{T}\mathcal{E}Qh+v^{T}Q^{-1}(\mathcal{E}^{2}+\mathcal{E}\mathcal{E}^{T}+\mathcal{E}^{2T})Q^{-T}v$
$dc_{II}(P;v,h)=h^{T}(\mathcal{E}^{T}P+P\mathcal{E})h-v^{T}(P^{-1}\mathcal{E}^{T}+\mathcal{E}P^{-1})v$
$\qquad\quad+h^{T}\mathcal{E}^{T}P\mathcal{E}h+v^{T}(\mathcal{E}^{2}P^{-1}+\mathcal{E}P^{-1}\mathcal{E}^{T}+P^{-1}\mathcal{E}^{2T})v$ (32)
where we have suppressed higher order terms $\mathcal{O}(\mathcal{E}^{3})$ to shorten our equations. The first choice in (32) not only looks more balanced and symmetric, but also has nicer properties than the second one as will be revealed in our following discussions. Thus, we mainly focus on the first choice.
One striking difference between fitting Hessian in the Lie groups and the manifold of SPD matrices is that now, criterion $c(P)$ could be strongly convex under certain mild conditions like $\lambda_{\min}(H)>0$. Clearly, $c(P)$ cannot be strictly convex in GL$(n,\mathbb{R})$ in the conventional sense as $Q$ and $UQ$ give the same value for $c(P)$ with any orthogonal matrix $U$. But, $c(P)$ could behave much better after properly separating out this rotation ambiguity. To this end, we need to turn to the polar decomposition [24] $A=U\Gamma$ to define a proper manifold that is irrelevant of such rotations, where $A\in\mathbb{R}^{n\times n}$ is a square matrix, $U$ is an orthogonal matrix, and $\Gamma\succeq 0$ is semi-positive definite. With the polar decomposition, let us introduce the following equivalence relation.
Equivalence relation $R_{\rm polar}$: For $Q_{1}$ and $Q_{2}$ in $\mathbb{R}^{n\times n}$, we say that $Q_{1}\sim Q_{2}$ if and only if their polar decompositions share the same $\Gamma$, i.e., $Q_{1}=U_{1}\Gamma$ and $Q_{2}=U_{2}\Gamma$ for certain orthogonal matrices $U_{1}$ and $U_{2}$.
It is not difficult to show that $R_{\rm polar}$ indeed defines an equivalence relation by showing that it is reflexive, symmetric and transitive. Then, we can define the quotient set GL$(n,\mathbb{R})/R_{\rm polar}$. Note that since the Lie group $O(n)$ has dimension $0.5n(n-1)$, elements from this quotient set have dimension $n^{2}-0.5n(n-1)=0.5n(n+1)$. Indeed, its elements can be represented as SPD matrices after discarding the rotation ambiguity. Hence, the SPD manifold can be defined as a quotient manifold in this way.
Proposition 5: Criterion $c(P)$ is strongly convex in the quotient set GL$(n,\mathbb{R})/R_{\rm polar}$ with the local coordinate system induced by choice I in (30) when $\lambda_{\min}(H)\geq\lambda_{0}>0$.
Proof: First, let us briefly show that with $A\succ 0$ and eigenvalue decomposition $A=UDU^{T}$, $aA+bA^{-1}-2\sqrt{ab}I=U(aD+bD^{-1}-2\sqrt{ab})U^{T}\succeq 0$, where $a$ and $b$ are two positive numbers.
Second, we show that it is sufficient to consider a symmetric $\mathcal{E}$ in the quotient set GL$(n,\mathbb{R})/R_{\rm polar}$. Starting from $Q$, let $Q_{1}=Q+\mathcal{E}_{1}Q$ be an arbitrary neighbor of $Q$ in the quotient set GL$(n,\mathbb{R})/R_{\rm polar}$ with $\mathcal{E}_{1}$ small enough such that $\|\mathcal{E}_{1}\|_{2}<1$. The condition, $\|\mathcal{E}_{1}\|_{2}<1$, is to ensure that $Q_{1}$ is invertible. Let the polar decomposition of $Q_{1}$ and $I+\mathcal{E}_{1}$ be
$Q_{1}=U_{1}\Gamma_{1},\qquad I+\mathcal{E}_{1}=U_{2}\Gamma_{2}$
respectively. Then, we have,
$\Gamma_{2}Q=U_{2}^{T}(I+\mathcal{E}_{1})Q=U_{2}^{T}Q_{1}=(U_{2}^{T}U_{1})\Gamma_{1}$
Thus, by letting $\mathcal{E}_{2}=\Gamma_{2}-I$, we have,
$(I+\mathcal{E}_{2})Q=(U_{2}^{T}U_{1})\Gamma_{1}\sim U_{1}\Gamma_{1}=(I+\mathcal{E}_{1})Q$ (33)
Since $\Gamma_{2}$ is symmetric, $\mathcal{E}_{2}=\Gamma_{2}-I$ is symmetric as well. Thus, (33) suggests that it is enough to only consider a symmetric $\mathcal{E}$ in the quotient set GL$(n,\mathbb{R})/R_{\rm polar}$. This is not surprising as a symmetric $\mathcal{E}$ just has dimension $0.5n(n+1)$.
The last step is to check the lower bound of the second order terms of $dc_{I}(P)$ in (32). Now, without loss of generality, let us assume that $\mathcal{E}=\mathcal{E}^{T}$ due to (33). Starting from (32), a lower bound of the expectation of the second order terms of $dc_{I}(P;v,h)$ with choice $dQ=\mathcal{E}Q$ is given by
$E_{v}\left[h^{T}Q^{T}\mathcal{E}^{T}\mathcal{E}Qh+v^{T}Q^{-1}(\mathcal{E}^{2}+\mathcal{E}\mathcal{E}^{T}+\mathcal{E}^{2T})Q^{-T}v\right]$
$=$ ${\rm tr}\left(\mathcal{E}^{2}QH^{2}Q^{T}+3\mathcal{E}^{2}Q^{-T}Q^{-1}\right)$
$\geq$ ${\rm tr}\left(\left(\lambda_{\min}(H)\right)^{2}\mathcal{E}^{2}QQ^{T}+3\mathcal{E}^{2}\left(QQ^{T}\right)^{-1}\right)$
$\geq$ ${\rm tr}\left(\mathcal{E}^{2}\left(\lambda_{0}^{2}QQ^{T}+3(QQ^{T})^{-1}\right)\right)$
$\geq$ $2\sqrt{3}\lambda_{0}\,{\rm tr}\left(\mathcal{E}^{2}\right)$
$=$ $2\sqrt{3}\lambda_{0}\left({\rm vec}(\mathcal{E})\right)^{T}{\rm vec}(\mathcal{E})$
$=$ $2\sqrt{3}\lambda_{0}\left({\rm vech}(\mathcal{E})\right)^{T}S^{T}S\,{\rm vech}(\mathcal{E})$
$=$ $\left({\rm vech}(\mathcal{E})\right)^{T}\left(2\sqrt{3}\lambda_{0}\,S^{T}S\right){\rm vech}(\mathcal{E})$
where we have used expectations $E_{v}[hh^{T}]=H^{2}$ and $E_{v}[vv^{T}]=I$ from the first to the second line, $aA+bA^{-1}\succ 2\sqrt{ab}I$ for $A\succ 0$ to arrive the fifth line from the fourth one, and definition ${\rm vec}(A)=S{\rm vech}(A)$ for a symmetric matrix $A$ to arrive the last line. Thus, the Hessian of $c_{I}(P)$ in the new coordinate is lower bounded by $2\sqrt{3}\lambda_{0}\,S^{T}S\succ 2\sqrt{3}\lambda_{0}I$ since $S^{T}S\succ I$. $\square$
We briefly present a few more statements regarding the convexity of the Hessian fitting problem in Lie groups without formal proofs. Clearly, $c(P)$ is strongly convex in any connected group of diagonal matrices when $\lambda_{\min}(H)\geq\lambda_{0}>0$. With group $O(n)$, from (32), it is not difficult to show that every point is a saddle point since $\mathcal{E}$$=-\mathcal{E}^{T}$. The second choice, $c_{II}(P)$, is generally not convex even with constraint $\mathcal{E}=\mathcal{E}^{T}$. Empirical results also have confirmed that Hessian fitting with choice II converges significantly slower than that with choice I. The first choice, $c_{I}(P)$, is generally not convex in the group of triangular matrices, although previous results have shown that Hessian fitting with this group does work well [17]. An explanation is given in the next subsection. Lastly, since we do not care about the rotation ambiguity for preconditioning purpose, Proposition 5 suggests that $c(P)$ is “virtually” strongly convex in the group GL$(n,\mathbb{R})$ with the local coordinate system induced by choice I in (30) when $\lambda_{\min}(H)\geq\lambda_{0}>0$.
III-B The Two Basic Lie Group Hessian Fitting Methods
III-B1 Fitting in Group GL$(n,\mathbb{R})$
As our Hessian fitting criterion is “virtually” strongly convex in this group, we can readily follow the conventional convex optimization theory to design proper Hessian fitting methods [1]. If we could find an upper bound for the eigenvalues of the Hessian of $c(P)$, say $L$, we know that GD or SGD with step size $1/L$ converges linearly to the optimal solution $H^{-1}$ at least with rate $1-2\sqrt{3}\lambda_{0}/L$, where $2\sqrt{3}\lambda_{0}$ is the lower bound of the eigenvalues of the Hessian of $c(P)$ derived from the proof of Proposition 5. From (32), choice I, we know that the sum of the second order terms of $\mathcal{E}$ is upper bounded by
$\left|h^{T}Q^{T}\mathcal{E}^{T}\mathcal{E}Qh+v^{T}Q^{-1}(\mathcal{E}^{2}+\mathcal{E}\mathcal{E}^{T}+\mathcal{E}^{2T})Q^{-T}v\right|$
$\leq$ $\left|h^{T}Q^{T}\mathcal{E}^{T}\mathcal{E}Qh\right|+\left|v^{T}Q^{-1}\mathcal{E}^{2}Q^{-T}v\right|$
$+\left|v^{T}Q^{-1}\mathcal{E}\mathcal{E}^{T}Q^{-T}v\right|+\left|v^{T}Q^{-1}\mathcal{E}^{2T}Q^{-T}v\right|$
$\leq$ $\left(\|Qh\|^{2}+3\|Q^{-T}v\|^{2}\right)\|\mathcal{E}\|^{2}$ (34)
where $\|\mathcal{E}\|$ can be any matrix norm for $\mathcal{E}$. Thus, from (34), we see that a rough estimate of $L$ could be
$L_{t}\approx\max_{1\leq i\leq t}\left(\|Q_{i}h_{i}\|^{2}+3\|Q_{i}^{-T}v_{i}\|^{2}\right)$ (35)
At the same time, from the linear terms in (32), choice I, we see that the gradient in group GL$(n,\mathbb{R})$ is
$\nabla_{Q}=2\left(Qhh^{T}Q^{T}-Q^{-T}vv^{T}Q^{-1}\right)$ (36)
which also is the gradient in the quotient set GL$(n,\mathbb{R})/R_{\rm polar}$ since it is already symmetric. Combining (34) and (36) gives the following first baseline Lie group SGD Hessian fitting method,
$Q_{t+1}=Q_{t}-\mu\frac{Q_{t}h_{t}h_{t}^{T}Q_{t}^{T}-Q_{t}^{-T}v_{t}v_{t}^{T}Q_{t}^{-1}}{\|Q_{t}h_{t}\|^{2}+\|Q_{t}^{-T}v_{t}\|^{2}}Q_{t}$ (37)
where $0<\mu\leq 2$ is a normalized step size. Note that in (37), we do not use the more conservative $L$ estimate given by (35). Numerical results suggest that (37) is stable enough.
A note on the matrix inverse in (37). Straightforward implementation of (37) needs to solve the linear system $Q^{T}x=v$ for calculating $Q^{-T}v$. Still, from (37), we see that $Q_{t+1}-Q_{t}$ is a matrix with rank two. Thus, we can recursively update the inverse of $Q$ with the Woodbury matrix identity [24] as done in BFGS. One concern is that the numerical errors may accumulate after repeatedly applying the Woodbury matrix identity many times. Empirical results have confirmed that this procedure is generally numerically stable with double and single precision arithmetic, but could be unstable with half precision arithmetic.
III-B2 Fitting in the Group of Triangular Matrices
We can simply take the triangular part of (36) to get the gradient for updating $Q$ in the group of triangular matrices as having done in [17]. But, in general $c(P)$ is not convex in this group. To take the advantage of Proposition 5, we can factorize out the rotation part of the updated $Q$ from (37) to have the following update formula for $Q$ in the group of upper triangular matrix,
$Q_{t+1}=\left[I-\mu\frac{Q_{t}h_{t}h_{t}^{T}Q_{t}^{T}-Q_{t}^{-T}v_{t}v_{t}^{T}Q_{t}^{-1}}{\|Q_{t}h_{t}\|^{2}+\|Q_{t}^{-T}v_{t}\|^{2}}\right]_{R}Q_{t}$ (38)
where $[A]_{R}$ only keeps the triangular part of the QR decomposition of a square matrix $A$. In this way, $Q_{t}$ is still in the quotient set GL$(n,\mathbb{R})/R_{\rm polar}$, but keeps to be a triangular matrix as long as $Q_{0}$ is. Note that only two rank-1 QR decomposition updates are required to obtain the QR decomposition of the matrix inside the operator $[\cdot]_{R}$ in (38) due to its low-rank structure [24]. One pleasant by-product is that now, we can calculate $Q_{t}^{-T}v_{t}$ with backward substitution, which is cheap and empirically shown to be numerically stable even with half precision arithmetic.
Let us derive an approximation for $[A]_{R}$ to reveal the connection between (38) and the method in [17]. Let $\|\Delta\|\ll 1$ be a small symmetric matrix, and consider the following QR decomposition,
$I+\Delta$ $=(I+\Delta_{1})(I+\Delta_{2})$
$I$ $=(I+\Delta_{1})^{T}(I+\Delta_{1})$ (39)
where $I+\Delta_{1}$ is an orthogonal matrix and $\Delta_{2}$ is an upper triangular matrix such that the first equation of (39) gives the QR decomposition for $I+\Delta$. Note that $\Delta_{1}$ and $\Delta_{2}$ are small matrices as well since QR decomposition is numerically stable. After expanding (39) and ignoring the second order terms $\Delta_{1}^{T}\Delta_{1}$ and $\Delta_{1}\Delta_{2}$, we can solve for $\Delta_{1}$ and $\Delta_{2}$ jointly from the following equations,
$\Delta_{1}+\Delta_{2}\approx\Delta,\qquad\Delta_{1}+\Delta_{1}^{T}\approx 0$
which can be shown to be,
$\Delta_{1}$ $\approx{\rm tril}(\Delta,-1)-{\rm triu}(\Delta,1)$
$\Delta_{2}$ $\approx{\rm triu}(\Delta)+{\rm triu}(\Delta,1)$
where ${\rm tril}(A,-1)$ and ${\rm triu}(A,1)$ exclude the diagonal elements of ${\rm tril}(A)$ and ${\rm triu}(A)$, respectively. Hence, we obtain the following approximation for $[A]_{R}$,
$\left[I+\Delta\right]_{R}\approx I+{\rm triu}(\Delta)+{\rm triu}(\Delta,1),\qquad{\rm for}\,\|\Delta\|\ll 1$ (40)
In this way, we see that the gradient in the group of triangular matrices still points to an ascent direction for updating $Q$ in the group GL$(n,\mathbb{R})$, although not the steepest one. It is safe to take the approximation (40) for the $[\cdot]_{R}$ operation in (38) for small update step sizes, say $\mu\leq 0.1$. Numerical results have confirmed that after separating out this rotation ambiguity, update (38) converges noticeably faster than the one given in [17]. Yet, for the case of large step sizes, the approximation in (40) loses precisions. Then, it is better to simply let $\left[I+\Delta\right]_{R}\approx I+{\rm triu}(\Delta)$ for $10\ll\|\Delta\|<$ 1.
III-B3 Newton’s Methods
Lastly, we want to briefly mention Newton’s methods for preconditioner fitting with Lie groups, which can be derived from (32) by solving for the optimal $\mathcal{E}$ that reduces $c(P)$ the most. But, they turn out to be significantly more complicated than the form in (15), and seem to be of little practical value. Nevertheless, the Hessian for the case considered in Proposition 5 still has the following neat form,
$S^{T}\left[I\otimes(QH^{2}Q^{T}+3Q^{-T}Q^{-1})\right]S$
III-C Sparse Lie Groups for Large Scale Hessian Fittings
The methods in (37) and (38) are only good for small to middle scale problems with roughly $n\leq 10^{5}$ parameters to optimize. For problems with millions to billions of parameters to learn, we need significantly sparser Hessian estimators.
III-C1 The Group of Diagonal Matrices
Let $Q={\rm diag}(q)$, where $q$ is a vector without zero element. Note that $\|A\|_{2}=\max_{1\leq i\leq n}|a_{i,i}|$ when $A$ is diagonal. It is trivial to following the same procedures leading to (37) to derive the following SGD method for updating $Q$ with this simple group,
$q_{t+1}=q_{t}-\mu\frac{\left(q_{t}h_{t}\right)^{2}-\left(v_{t}/q_{t}\right)^{2}}{\max\left(\left(q_{t}h_{t}\right)^{2}+\left(v_{t}/q_{t}\right)^{2}\right)}q_{t}$ (41)
where /// denote element-wise division if its two operands are vectors.
III-C2 Decomposable Groups for Matrix Parameters
A $Q$ that can be decomposed into form $Q=\oplus_{i=1}^{I}\left(Q_{i,2}\otimes Q_{i,1}\right)$ can be useful for neural network training since those models mainly consist of a list of modules using matrices or tensors as trainable parameters, see [18] for a detailed treatment. Specifically, let $\Theta$ be a matrix parameter, and $y=\Theta x$ be the output of a certain layer with input feature vector $x$, where the last element $x$ can be optionally set to $1$ such that this is a general affine transform. Then, the PSGD iteration in (1) with $Q=Q_{2}\otimes Q_{1}$ reduces to an SGD update in a $\Theta$-dependent new coordinate defined by $\Theta^{\prime}=Q_{1}^{-T}\Theta Q_{2}^{-1}$ with affine transform $y^{\prime}=\Theta^{\prime}x^{\prime}$, where $y^{\prime}=Q_{1}^{-T}y$ and $x^{\prime}=Q_{2}x$ [18]. Hence, we call this type of preconditioner the affine one. Let us derive an improved implementation for fitting such preconditioners in the light of Proposition 5.
Without loss of generality, we only consider the case of $Q=Q_{2}\otimes Q_{1}$. First, let us freeze $Q_{1}$ and only consider the update of $Q_{2}$. By Proposition 5, we should take $dQ_{2}=\mathcal{E}_{2}Q_{2}$, and thus,
$dQ=\left(\mathcal{E}_{2}Q_{2}\right)\otimes Q_{1}=\left(\mathcal{E}_{2}\otimes I\right)\left(Q_{2}\otimes Q_{1}\right)=\left(\mathcal{E}_{2}\otimes I\right)Q$
Then, by replacing the $\mathcal{E}$ in (32) with $\left(\mathcal{E}_{2}\otimes I\right)$, we obtain the second order approximation for $c(P+dP)-c(P)$. Also, noting that $\|\mathcal{E}_{2}\otimes I\|_{2}=\|\mathcal{E}_{2}\|_{2}$, we can follow the procedures leading to (38) to bound the Hessian with respect to $\mathcal{E}_{2}$ to derive a step size normalized update rule for $Q_{2}$. In the same way, we can freeze $Q_{2}$ and have $dQ=\left(I\otimes\mathcal{E}_{1}\right)Q$ by taking $dQ_{1}=\mathcal{E}_{1}Q_{1}$. Again, noting that $\|I\otimes\mathcal{E}_{1}\|_{2}=\|\mathcal{E}_{1}\|_{2}$, we can derive another step size normalized update rule for $Q_{1}$. We skip these tedious details, and directly give the SGD update for $Q$ as below,
$A_{t}$ $=Q_{1,t}{\rm uvec}(h_{t})Q_{2,t}^{T},\;B_{t}=Q_{2,t}^{-T}\left[{\rm uvec}(v_{t})\right]^{T}Q_{1,t}^{-1}$
$Q_{1,t+1}$ $=\left[I-\mu\,\frac{A_{t}A_{t}^{T}-B_{t}^{T}B_{t}}{\|A_{t}A_{t}^{T}+B_{t}^{T}B_{t}\|_{2}}\right]_{R}Q_{1,t}$
$Q_{2,t+1}$ $=\left[I-\mu\,\frac{A_{t}^{T}A_{t}-B_{t}B_{t}^{T}}{\|A_{t}^{T}A_{t}+B_{t}B_{t}^{T}\|_{2}}\right]_{R}Q_{2,t}$ (42)
where ${\rm uvec}(v_{t})$ and ${\rm uvec}(h_{t})$ recover vectors $v_{t}$ and $h_{t}$ to their original matrix forms, respectively.
The set of equations in (42) looks almost the same as that in [17] in Section V.B for updating $Q$ except that here, we have normalized the gradients with spectral norms and removed the rotation ambiguity by using the operator $[\cdot]_{R}$ approximation in (40). Again, in practice, the matrix inverses in (42) should be replaced with backward substitutions, which are as cheap and numerically stable as matrix multiplications. Also, note that (42) reduces to (38) by choosing $Q_{2}=1$ when $\Theta\in\mathbb{R}^{n\times 1}$.
Note that the KFAC [5] and Shampoo [6, 7, 8, 9] like preconditioners are quite different from ours even though they all share the same Kronecker product forms. The KFAC and Shampoo preconditioners can only approximate the inverse of Hessians even when they can be factorized into the assumed Kronecker product forms, while our affine one could converge to the exact solution given the correct factorization form.
III-C3 Decomposable Groups for Tensor Parameters
Similar to the above affine preconditioners, a $Q$ that can be decomposed into the more general form $Q=\oplus_{i=1}^{I}\left(\otimes_{j=1}^{J_{i}}Q_{i,j}\right)$ is useful for preconditioning the gradients of a list of higher order tensors. To simplify our equations, let us adopt the Einstein summation notation, only consider the preconditioning of one tensor, and suppress the time index. For a tensor $\mathcal{T}$, we precondition its gradient $\mathcal{G}$ with the following multilinear mapping,
$\mathcal{G}^{\rm preconditioned}_{ab\ldots}=(Q_{1})_{a\alpha}(Q_{2})_{b\beta}\ldots\mathcal{G}_{\alpha\beta\ldots}$
Given pair $(v,h)$, we restore them to the same shape as $\mathcal{T}$, and prepare the following two tensors,
$\mathcal{A}_{ab\ldots}=$ $(Q_{1})_{a\alpha}(Q_{2})_{b\beta\ldots}({\rm uvec}(h))_{\alpha\beta\ldots}$
$\mathcal{B}_{ab\ldots}=$ $({\rm uvec}(v))_{\alpha\beta\ldots}(Q_{1}^{-1})_{\alpha a}(Q_{2}^{-1})_{\beta b\ldots}$
Then, taking $Q_{j}$ as an example, the preconditioner can be updated as
$(Q_{j})_{ac}\leftarrow\left[I_{ab}-\mu\frac{(C(\mathcal{A},j))_{ab}-(C(\mathcal{B},j))_{ab}}{\left\|C(\mathcal{A},j)+C(\mathcal{B},j)\right\|_{2}}\right]_{R}(Q_{j})_{bc}$
where $C(\mathcal{A},j)$ is the matrix obtained by ${\rm einsum}(\mathcal{A},\mathcal{A})$ that only keeps the $j$th indix of its two input terms and contracts away all the other indices. Matrix $C(\mathcal{B},j)$ is calculated in the same way. Please refer to our supplementary materials [26] for detailed derivations.
III-C4 Low-Rank Approximation (LRA) Preconditioner
Similar to the L-BFGS method, we have developed a LRA preconditioner in [19]. Let us revisit it and refine its implementations. We can have two choices for the definition of $Q$,
${\rm choice\,I:}\quad Q$ $=(I+UV^{T}){\rm diag}(d)$
${\rm choice\,II:}\quad Q$ $={\rm diag}(d)(I+UV^{T})$ (43)
where $U$ and $V\in\mathbb{R}^{n\times r}$ with $0\leq r\ll n$, and vector $d$ has no zero element such that ${\rm diag}(d)$ is in the group of diagonal matrices. Both choices of (43) have the same representation capacity for nonsingular matrices with form ${\rm diag}(d^{\prime})+U^{\prime}V^{\prime T}$. Still, the first choice is easier to work with when we are to bound the Hessians with respect to $U$ and $V$.
The key to work with the LRA preconditioner is to notice the Lie group structures of the $Q$ in (43). Clearly, ${\rm diag}(d)$ is already in the group of diagonal matrices. Noting that
$e^{\varepsilon G_{1}V^{T}}-I=\left(\sum_{i\geq 1}\frac{\varepsilon^{i}(G_{1}V^{T})^{i-1}G_{1}}{i!}\right)V^{T}=\mathcal{E}_{1}V^{T}$ (44)
readers familiar with Lie group can see that set $\{I+UV^{T}|U\in\mathbb{R}^{n\times r},\,\det(I+V^{T}U)\neq 0,V{\rm\,is\,fixed}\}$ forms a Lie group. Similarly, with a given $U$, set $\{I+UV^{T}|V\in\mathbb{R}^{n\times r},\,\det(I+V^{T}U)\neq 0,U{\rm\,is\,fixed}\}$ forms another Lie group. We can exploit these Lie group structures to update $U$, $V$ and $d$ efficiently. For example, to update $U$, from (44), we can let $dQ$ be
$dQ=e^{\varepsilon G_{1}V^{T}}Q-Q=\mathcal{E}_{1}V^{T}Q$
and then simply replace the $\mathcal{E}$ in (32) with $\mathcal{E}_{1}V^{T}$ to obtain the second order approximation of $c(P+dP)-c(P)$. Similarly, to update $V$, we switch to the other group and let $dQ$ be
$dQ=e^{\varepsilon UG_{2}^{T}}Q-Q=U\mathcal{E}_{2}^{T}Q$
We omit the lengthy derivations here and only present the final update rules for $U$, $V$ and $d$ as below,
$a_{t}=$ $Q_{t}h_{t},\quad b_{t}=Q_{t}^{-T}v_{t}$
$d_{t+1}=$ $\left(1-\mu\frac{(Q_{t}^{T}a_{t})h_{t}-v_{t}(Q_{t}^{-1}b_{t})}{\max\sqrt{\left((Q_{t}^{T}a_{t})^{2}+v_{t}^{2}\right)\left(h_{t}^{2}+(Q_{t}^{-1}b_{t})^{2}\right)}}\right)d_{t}$
$U_{t+1}=$ $U_{t}-\mu\frac{(a_{t}a_{t}^{T}-b_{t}b_{t}^{T})V_{t}(I+V_{t}^{T}U_{t})}{\|a_{t}\|\|V_{t}V_{t}^{T}a_{t}\|+\|b_{t}\|\|V_{t}V_{t}^{T}b_{t}\|}$
$V_{t+1}=$ $V_{t}-\mu\frac{(I+V_{t}U_{t}^{T})(a_{t}a_{t}^{T}-b_{t}b_{t}^{T})U_{t}}{\|a_{t}\|\|U_{t}U_{t}^{T}a_{t}\|+\|b_{t}\|\|U_{t}U_{t}^{T}b_{t}\|}$ (45)
Interested readers can refer to the supplementary materials of [19] for further details. The only difference is that in (45), we have normalized the step sizes with information from the second order derivatives.
A few notes before ending the LRA preconditioner discussions. The matrix inverse in (45) can be cheaply calculated with the Woodbury matrix identity [24] since typically $0\leq r\ll n$. We usually update either $U$ and $V$, not both, within a single update step as set $\{I+UV^{T}|U,V\in\mathbb{R}^{n\times r}\,{\rm and}\,\det(I+V^{T}U)\neq 0\}$ no longer forms a Lie group. Our LRA preconditioner is quite different from the one in L-BFGS. The L-BFGS is basically a limited number of steps of expansion of the BFGS iterations, and has no memories of Hessian-vector products beyond $r$ steps. Clearly, (45) is not a simple expansion of (37) up to $r$ steps. It keeps track of the best approximation of $H^{-1}$ with a diagonal plus a low-rank matrix. Also, it is not difficult to check that with $Q={\rm diag}(q)$, we have,
$\left((Q_{t}^{T}a_{t})^{2}+v_{t}^{2}\right)\left(h_{t}^{2}+(Q_{t}^{-1}b_{t})^{2}\right)=\left(\left(q_{t}h_{t}\right)^{2}+\left(v_{t}/q_{t}\right)^{2}\right)^{2}$
which suggests that (45) reduces to (41) when $r=0$.
III-C5 More Sparse Lie Groups for Preconditioner Fittings
It is not possible to enumerate all the potentially useful sparse Lie groups for preconditioner fittings. One systematic way to create such sparse Lie groups is to construct the so-called matrix-free preconditioners from any subgroups of the (discrete) permutation group with order $n$ [19]. Specifically, the diagonal preconditioner is the simplest case from this family, induced by the identity element, $\{e\}$, from the permutation group. For an even $n$, $\{e,s_{n/2}\}$ induces the butterfly matrices which also form a Lie group, where $s_{n/2}$ denotes the $n/2$-circular-shift operator [19]. All these groups can be reduced to the direct sum of a series of smaller ones.
Another interesting and nontrivial Lie group is the set of invertible matrices having form [18]
$Q=\begin{bmatrix}{\rm bdiag}(A_{1},\ldots,A_{i})&B\\ 0&C\end{bmatrix}$ (46)
where ${\rm bdiag}(A_{1},\ldots,A_{i})$ is a block diagonal matrix, $C$ and all the blocks $A_{j}$, $1\leq j\leq i$, are invertible, and $B$ is a matrix with proper size. It is called the normalization Lie group in [18] since as a special affine transform, the feature normalization matrices do have this form. For example, in the batch [15] or layer normalizations [16], a feature vector $x$ is partitioned into $i$ groups as $x=[x_{1};\ldots;x_{i};1]$ and normalized as,
$\begin{bmatrix}x_{1}^{\prime}\\ \vdots\\ x_{i}^{\prime}\\ 1\end{bmatrix}=\begin{bmatrix}\frac{1}{a_{1}}I&&&-\frac{b_{1}}{a_{1}}e\\ &\ddots&&\vdots\\ &&\frac{1}{a_{i}}I&-\frac{b_{i}}{a_{i}}e\\ &&&1\end{bmatrix}\begin{bmatrix}x_{1}\\ \vdots\\ x_{i}\\ 1\end{bmatrix}$ (47)
where $e$ is a vector of ones, and $a_{j}$ and $b_{j}$, $1\leq j\leq i$, are the shared standard derivation and mean of the $j$th group, respectively. It is not difficult to check that matrices with the form in (47) do form a Lie group. Due to the importance of feature normalizations in neural network learning [15, 16], this simple Lie group is useful by providing a tool for fitting such preconditioners in an elegant and principled way.
III-D Summary of Lie Group Hessian Fitting Methods
Table III summarizes the information of the above five basic types of Lie group preconditioners. With the diagonal preconditioner as the base line and assuming that $\theta={\rm vec}(\Theta)$ with $\Theta\in\mathbb{R}^{m\times m}$, we see that the two dense preconditioner fitting methods have quadratic or high complexities. The affine and LRA preconditioners are two promising choices. Between these two preconditioners, the affine one requires more computations, while the LRA one with $r\ll m$ needs more memory spaces. It is possible to mix these basic building blocks to derive richer forms of Lie group preconditioners, e.g., the scaling-normalization preconditioner in [18] which only has a sublinear space complexity, i.e., $\mathcal{O}(m)$, although its computational complexity still is $\mathcal{O}(m^{2})$.
TABLE III: Lie groups for Hessian fittings. Assuming that $\theta={\rm vec}(\Theta)$ with $\Theta\in\mathbb{R}^{m\times m}$ for the storage and computation numbers.
$Q$ $Q_{t+1}\leftarrow Q_{t}$ Storages Computations
GL$(n,\mathbb{R})$ (37) $\mathcal{O}(m^{4})$ $\mathcal{O}(m^{4})$
Triangular (38) $\mathcal{O}(m^{4})$ $\mathcal{O}(m^{6})$
Diagonal (41) $\mathcal{O}(m^{2})$ $\mathcal{O}(m^{2})$
$Q_{2}\otimes Q_{1}$ (42) $\mathcal{O}(m^{2})$ $\mathcal{O}(m^{3})$
$(I+UV^{T}){\rm diag}(d)$ (45) $\mathcal{O}(rm^{2})$ $\mathcal{O}(rm^{2})$
Fig. 2 shows a few typical Hessian fitting convergence curves for the PSGD Hessian fitting methods with (37) and (38) using default step size $\mu=1$, the closed-form solution (18) by clipping the maximum moving average factor to $0.999$, and the classic BFGS formula [1]. We have tested the following three cases.
III-D1 Clean Hessian-Vector Products
In this case, we set the ground truth $H^{\prime}$ to $H$. All methods except for the closed-form one can converge to the ground truth solution up to errors only bounded by the machine precision. The PSGD style with (37) and BFGS have higher error plateaus than the PSGD style with (38) due to the accumulated numerical errors caused by repeated usage of the Woodbury matrix identity [24] for recursive matrix inverse updating.
III-D2 Noisy Hessian-Vector Products
We have tested a noisy model of (3) with $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}I)$ and $\sigma_{\epsilon}=0.01$. Hence, we set the ground truth to $H^{\prime}=(H^{2}+\sigma_{\epsilon}^{2}I)^{0.5}$ due to (5). We see that the BFGS is very sensitive to noises, and diverges easily. The two PSGD methods with reduced step size $\mu=0.1$ converge slower than the closed-form solution initially, but reach lower fitting errors eventually. Note that now there always are certain steady state preconditioner fitting errors with PSGD since the gradient in (36) cannot settle down to zero when $\sigma_{\epsilon}>0$. Yet, the preconditioner fitting problem is still strongly convex. One can show this by following the proof of Proposition 5 and replacing the $H$ there with $(H^{2}+\sigma_{\epsilon}^{2}I)^{0.5}$. If necessary, the preconditioner fitting errors of PSGD can be sufficiently small given enough iterations to anneal down the step sizes.
III-D3 Time-Varying Hessians
The Hessians are from a time varying process defined by $H_{t+1}=H_{t}+uu^{T}$, where $u_{i}\sim\mathcal{U}(0,1)$ for $1\leq i\leq n$, and $H_{0}$ is a matrix with all elements being $1/4$. The ground truth $H^{\prime}=H_{t}$ is also time-varying. The BFGS diverges first as the Hessians change too fast, and only locks onto the target $H_{t}^{-1}$ after $t>1000$. The closed-form solution has difficulty in tracking a time-varying Hessian due to its sublinear convergence. The two PSGD style Hessian fitting methods lock onto the target quicker than BFGS, and also do not show any sign of divergence before convergence.
Figure 2: (a) Typical convergence curves when $H$ is a $50\times 50$ matrix with $H_{i,i}=1$ and $H_{i,j}=0.5$ for $|i-j|=1$. Its eigenvalues are bounded in range $(0,2)$. (b) The same $H$ as in (a), but with a noisy model of (3), where $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}I)$, and $\sigma_{\epsilon}=0.01$. (c) Time varying Hessians defined by process $H_{t+1}=H_{t}+uu^{T}$, where $u_{i}\sim\mathcal{U}(0,1)$ for $501\leq i\leq n=$ 50, and $H_{0}$ is a matrix with all elements being $1/4$.
IV More Experimental Results and Discussions
So far we have detailed the behaviors of a wide range of Hessian fitting methods, see the summaries in Table II and III, and Fig. 1 and Fig. 2. Here, we present a few more experimental results to demonstrate their typical behaviors in practice. Interested readers can refer to our reports in [21] for significantly more extensive comparison results among these Hessian fitting methods and many more optimizers on a wide range of machine learning benchmarks. Focus here is on the general properties of different stochastic Hessian fitting methods. In general, we do observe that accurate Hessian estimation typically leads to faster convergences and better generalizations. In this section, the SGD only refers to the one for learning the parameter $\theta$, not the one for Hessian fitting. The PSGD implementations refer to the improved versions in this paper, not the original ones from [17, 18, 19]. For PSGD, the default step size for preconditioner fitting is typically set to $0.1$ and $1.0$ for stochastic and convex optimizations, respectively, if no divergence issue. The preconditioner is initialized to $P_{0}=\alpha I$. Given pair $(v_{0},h_{0})$, it is ready to find that the optimal $\alpha$ minimizing criterion (2) is $\alpha=\|v_{0}\|/\|h_{0}\|$. In this way, PSGD is almost a tuning-free optimizer.
IV-A Applications to Stochastic Optimizations
IV-A1 Accuracy Comparisons among Different Preconditioner Fitting Methods
For the minimization of NLL losses, the Fisher and empirical Fisher information matrices (FIM) are closely related to the Hessian and their inverses are commonly used as the preconditioners. In this experiment, we check the preconditioner estimation with pairs $(v,g)$ and $(v,h)$. From Table I, we see that for the closed-form solutions, $v$ is a dummy variable and does not shows up in the solutions for $P$ since $E_{v}[v^{T}P^{-1}v]={\rm tr}(P^{-1})$. For preconditioner fitting with PSGD using pair $(v,g)$, $v$ can be optionally kept as an auxiliary variable so that we can reuse the same math and code for fitting $P$ with pair $(v,h)$. Fig. 2 already illustrates the performance of the exact closed-form solution. Here, we only consider approximate closed-form solutions like KFAC [5] and Shampoo [7]. Specifically, with gradient $G={\rm uvec}(g)$ for matrix parameter $\Theta$, Shampoo uses closed-form solution
$\left(E[G^{T}G]\right)^{-1/p}\otimes\left(E[GG^{T}]\right)^{-1/p}$ (48)
as the preconditioner for $g$, where $p=2$ or $4$. Note that (48) is not an exact solution even when $E[gg^{T}]$ can be perfectly factorized into the Kronecker product of two matrices with the assumed shapes.
Fig. 3 shows the typical convergence curves for a few methods assuming that the Hessian is decomposable. The closed-form solution Shampoo [7] and its recently improved version, Combining AxeS PReconditioners (CASPR) [9], can only roughly approximate the true solution and perform the worst. The steady-state fitting errors of PSGD-Affine with pair $(v,g)$ cannot reduce to zero with a constant finite step size, which is $\mu=0.1$ here. Still, it significantly outperforms Shampoo like coarse approximations. Also, PSGD-Affine requires less memories than Shampoo as it consumes $g$ on the fly, while Shampoo needs to buffer the matrices $E[GG^{T}]$ and $E[G^{T}G]$, as suggested by (48). PSGD-Affine with pair $(v,h)$ performs the best, and converges to $H^{-1}$ within errors only bounded by the machine precisions. This is expected as the pair $(v,h)$ carries more information on $H$ than $g$ alone. In practice, we also observe that Shampoo like methods are biased and may need to be grafted onto Adam or RMSProp to work properly.
Figure 3: Stochastic gradients and Hessian-vector products for preconditioner fitting, where $H\in\mathbb{R}^{100\times 100}$ is decomposable as $H=H^{\prime}\otimes H^{\prime}$ with $H^{\prime}_{i,i}=1$ and $H^{\prime}_{i,j}=0.5$ for $|i-j|=1$. Note that the convergence curves of Shampoo and CASPR are too close to be distinguishable.
IV-A2 Comparison with Shampoo
This example compares PSGD with Shampoo [7] on the LeNet5 [25] convolutional network for handwritten digits recognitions. Fig. 4 shows the typical convergence curves of SGD, Shampoo and PSGD-Affine with batch size $64$ and proper tunings for each method. Shampoo has higher time complexity than PSGD-Affine as the most expensive operations of Shampoo, i.e., the eigenvalue decompositions (EVD) for calculating the inverse of matrix $p$th root in (48), are far less efficient than the most expensive operations of PSGD-Affine, i.e., the backward substitutions in (42). The backward substitution is roughly $20$ to $30$ times faster than EVD for matrices with dimensions up to $10^{3}$ or larger in our test system equipped with an Nvidia RTX 2080 graphic card. Regarding the test classification error rates, the mean numbers averaged over ten runs are $0.95\%$, $0.8\%$ and $0.7\%$ for SGD, Shampoo and PSGD-Affine, respectively.
Figure 4: (a) Train loss versus epochs for the LeNet5 benchmark. (b) The same as (a) except that the x-axis shows the wall-time running on an Nvidia RTX 2080 graphic card.
IV-A3 Comparison with L-BFGS
We consider the same task as the above experiment, but replace LeNet5 with a logistic regression model using features $\{1,x_{i,j},x_{i_{1},j_{1}}x_{i_{2},j_{2}}\}$, where $x_{i,j}$, $1\leq i,j\leq 28^{2}$, is the pixel value at location $(i,j)$. This results in a large-scale logistic regression problem with over six millions coefficients to fit. L-BFGS is the selected method for this type of convex optimization problem [1]. Since BFGS is sensitive to gradient noises as shown in Fig. 2, we set batch size to the maximum allowable number on our hardware, $500$. Still, L-BFGS has a good chance, roughly $1/3$, to diverge. Fig. 5 shows the typical convergence curves of SGD, L-BFGS and PSGD-LRA with proper hyper-parameter tunings. From Fig. 5, we see that the line search step of L-BFGS not only makes it have a higher time complexity than PSGD-LRA, but also a higher regression loss plateau. Regarding generalization performance, mean test classification error rates over ten runs are about $2\%$ for L-BFGS and PSGD-LRA, and $2.3\%$ for SGD.
Figure 5: (a) Train loss versus epochs for the logistic regression benchmark. (b) The same as (a) except that the x-axis shows the wall-time running on an Nvidia RTX 2080 graphic card.
IV-B Applications to the Conventional Convex Optimizations
We would like to briefly address the traditional convex or mathematical optimizations [1] before ending this report. The results shown in Fig. 2 clearly suggest the potential of PSGD to such applications. What is interesting is that PSGD seems to be less vulnerable to saddle points than those off-the-shelf optimizers since it does not rely on line search. Let us consider a tensor rank decomposition problem defined by
$\min_{x,y,z}\sum_{i=1}^{I}\sum_{j=1}^{J}\sum_{k=1}^{K}\left(\tau_{i,j,k}-\sum_{r=1}^{R}x_{ri}y_{rj}z_{rk}\right)^{2}$ (49)
where $\tau$ is a given tensor, and parameters $x,y$ and $z$ are to be solved. The loss landscape is riddled with saddle points, e.g., $x=y=0$ and an arbitrary $z$. Fig. 6 shows a set of typical convergence curves from a few general purpose optimizers. The stairway like convergence curve of L-BFGS clearly indicates that line search has difficulty in escaping saddle points as the gradients around there are too small. PSGD works equally well for positive and non-positive definite Hessian matrices, and is far less vulnerable to saddle points.
Figure 6: Comparisons on the tensor rank decomposition problem of (49), where $R=10,I=20,J=50$, $K=100$, $\tau_{i,j,k}=\sum_{r=1}^{R}a_{ri}b_{rj}c_{rk}$, and all the elements of $a$, $b$ and $c$ are drawn from $\mathcal{N}(0,1)$. The XMat preconditioner is a simple matrix-free one proposed in [19].
Nevertheless, line search is an important step of convex optimizations for improved convergence. For example, let us consider the strongly convex loss $f(\theta)=\theta^{2}-4\sqrt{\theta}$ with $\theta\in\mathbb{R}^{+}$. Starting from initial guess $\theta=0$, Newton’s method will get stuck there since $\lim_{\theta\rightarrow 0}\frac{df(\theta)}{d\theta}/\frac{d^{2}f(\theta)}{d\theta^{2}}=-2\lim_{\theta\rightarrow 0}\theta=0$. GD or line search helps to push $\theta$ to the basin of attraction, only where Newton’s method converges quadratically. The purpose of the above experiment is to show that PSGD can operate without line search. This could bring new opportunities for enhancing the traditional BFGS like quasi-Newton optimizers by relaxing their line search conditions.
V Conclusion
In this paper, we have studied a wide range of stochastic Hessian fitting methods using stochastic gradients or Hessian-vector products as inputs. Exact closed-form solutions are available only for a few simple cases like a diagonal or a dense preconditioner form. They do converge sublinearly to the optimal solution, but may involve numerically challenging matrix operations like inverses and eigenvalue decompositions. The stochastic Hessian fitting criterion from the preconditioned stochastic gradient descent (PSGD) method is thoroughly analyzed in the Euclidean space, the manifold of symmetric positive definite (SPD) matrices and a variety of Lie groups. It is shown to be strongly convex under mild conditions for the Lie group GL$(n,\mathbb{R})$ with one form of local coordinate definition, see the Proposition 5 in Section III.A for details. This unique discovery facilitates our designs of several Lie group stochastic Hessian fitting methods with linear convergences, self-normalized step sizes, and different time and space complexities, see Table III in Section III.D for a summary. Practically, these Lie group stochastic Hessian fitting methods successfully avoid any numerically problematic operations like large matrix inverse or decompositions. Numerical results have confirmed our theoretical analyses, and demonstrated their efficiency and numerical reliability for real world stochastic optimizations.
Acknowledgment
I am deeply grateful to those pioneers, especially Omead Pooladzandi and Evan Walters, who have successfully applied PSGD to a wide range of supervised and reinforcement learning problems, revealed the great potential and elegancy of these Lie group based optimization methods for extremely large-scale problems, and shared with me their valuable feedback and insights.
Appendix A: Implementation Details
I have provided a basic Pytorch PSGD implementation at https://github.com/lixilinx/psgd_torch. Links to a few more specialized Jax and Pytorch implementations are provided there as well.
V-A Matrix Spectral Norm Estimation
The spectral norm is used to bound the Lipschitz constant of the PSGD preconditioner fitting criterion. Its estimation is crucial to the preconditioner update rule of PSGD. There are abundant of matrix norm estimators. But, we prefer some very computationally cheap and still tight enough ones.
Given a matrix $A\in\mathbb{R}^{n\times n}$, let us introduce notations
$\|A\|_{\max}$ $={\max}_{1\leq i,j\leq n}|a_{ij}|$
$\|A\|_{1}$ $={\max}_{1\leq j\leq n}\sum_{i=1}^{n}|a_{ij}|$
$\|A\|_{\infty}$ $={\max}_{1\leq i\leq n}\sum_{j=1}^{n}|a_{ij}|$
$\beta$ $=\sqrt{{\max}\left({\max}_{1\leq i\leq n}\sum_{j=1}^{n}a_{ij}^{2},\,\max_{1\leq j\leq n}\sum_{i=1}^{n}a_{ij}^{2}\right)}$
and give the following bounds for $\|A\|_{2}$ without proof,
$\frac{1}{\sqrt{r}}\|A\|_{F}$ $\leq\|A\|_{2}\leq\|A\|_{F}$
$\|A\|_{\max}$ $\leq\|A\|_{2}\leq n\|A\|_{\max}$
$\frac{1}{\sqrt{n}}\|A\|_{1}$ $\leq\|A\|_{2}\leq\sqrt{n}\|A\|_{1}$
$\frac{1}{\sqrt{n}}\|A\|_{\infty}$ $\leq\|A\|_{2}\leq\sqrt{n}\|A\|_{\infty}$
$\beta$ $\leq\|A\|_{2}\leq\sqrt{n}\beta$
where $r$ is the rank of $A$. Except when $r\ll n$, $\beta$ gives the tightest lower bound for $\|A\|_{2}$. Still, we can further tighten it by a few power iterations. Without loss of generality, we assume that $\beta$ equals the vector norm of the $i$th column of $A$, i.e., $a_{i}$. Then, with only two matrix-vector products, a significantly tighter bound is given by
$\|A\|_{2}\geq\frac{\|AA^{T}a_{i}\|}{\|A^{T}a_{i}\|}$
V-B Negative Log Likelihood Losses and Gradient Whitening
V-B1 Relationship between Hessian and FIM
Most machine learning optimization problems exclusively deal with the minimization of negative log likelihood (NLL) losses. Let $f(x;\theta)=\log p(x;\theta)$ be a differentiable logarithm probability density function (pdf) of random variable $x$ parametrized with $\theta$. Due to the relationship,
$E\left[\left(\frac{\partial f(x;\theta)}{\partial\theta}\right)\left(\frac{\partial f(x;\theta)}{\partial\theta}\right)^{T}\right]=-E\left[\frac{\partial^{2}f(x;\theta)}{\partial\theta^{T}\partial\theta}\right]$
we see that the Hessian, Fisher information matrix (FIM) and empirical FIM matrix all are the same when $z$ indeed follows the assumed form of distribution. Specifically, let $g$ be the gradient averaged over $B$ i.i.d. samples, and $m_{t}$ be the momentum updated as
$m_{t+1}=\beta m_{t}+(1-\beta)g_{t}$
where $10\leq\beta<$ 1. Then, we have
${\rm Hessian}={\rm FIM}=B\,E_{z}[gg^{T}]=B\frac{1+\beta}{1-\beta}E_{z}[mm^{T}]$ (50)
Hence, it is well motivated to use the gradient or momentum whitening preconditioner for NLL minimization. Note that the four terms in (50) are not necessarily the same when $z$ deviates from the assumed form of distribution.
Let us consider a simple example to clarify the above statements. Let $z_{i}$, $1\leq i\leq T$, be the given data points. We are to fit their distribution as a normal one with unknown $\mu$ and $\sigma^{2}$. Hence, the parameters are $\theta=(\mu,\sigma)$. The NLL is given by
$E_{z}\left[(z-\mu)^{2}/(2\sigma^{2})+\log\sigma\right]$
where $E_{z}[\cdot]$ dontes the sample average over $z$. The per-sample gradient and Hessian are given by
$g_{z}=$ $\begin{bmatrix}(\mu-z)/\sigma^{2}\\ 1/\sigma-(\mu-z)^{2}/\sigma^{3}\end{bmatrix}$
$H=$ $E_{z}\begin{bmatrix}1/\sigma^{2}&2(z-\mu)/\sigma^{3}\\ 2(z-\mu)/\sigma^{3}&3(\mu-z)^{2}/\sigma^{4}-1/\sigma^{2}\end{bmatrix}$
respectively. The FIM is ${\rm diag}(1/\sigma^{2},2/\sigma^{2})$, which depends on the assumed pdf, i.e., $\mathcal{N}(\mu,\sigma^{2})$, not on the actual distribution of $z$. The empirical Fisher, $E_{z}[g_{z}g_{z}^{T}]$, is not necessarily diagonal in general. The Hessian is even not necessarily positive definite, say for $\sigma\rightarrow\infty$. Nevertheless, they all reduce to the same form when $z\sim\mathcal{N}(\mu,\sigma^{2})$. Under the weaker condition of $E_{z}[g_{z}]=0$, we see that $H$ still reduces to the FIM, but the empirical FIM can be different from the true FIM if $z$ is not Gaussian. In general, we can show that $H$ and FIM are the same when $E_{z}[g_{z}]=0$ and $p(\cdot;\theta)$ is from the exponential family. Further discussions of these topics are beyond the scope of this report.
V-B2 On the Auxiliary Variable $v$
One more topic is on whether we should integrate out $v$ for updating the gradient whitening preconditioner as it appears to be a nuisance variable in (2). Actually, the term $v^{T}Pv$ can be regarded as a Hutchinson trace estimator as its expectation is ${\rm tr}(P)$. Since $E_{v\sim\mathcal{N}(0,I)}[Q^{-T}vv^{T}Q^{-1}]=Q^{-T}Q^{-1}$, (37) becomes
$Q_{t+1}=Q_{t}-\mu\frac{Q_{t}g_{t}g_{t}^{T}Q_{t}^{T}-Q_{t}^{-T}Q_{t}^{-1}}{\|Q_{t}g_{t}\|^{2}+{\rm tr}(Q_{t}^{-T}Q_{t}^{-1})}Q_{t}$ (51)
after integrating out $v$ and replacing $h$ with $g$. Note that for $Q\in\mathbb{R}^{n\times n}$, (37) only has computational complexity $\mathcal{O}(n^{2})$ while (51) has complexity $\mathcal{O}(n^{3})$. In general, we prefer to keep $v$ as an auxiliary variable to simplify the code and computations, although I have provided the implementations for both cases for the Kron product preconditioner.
V-C On the LRA Groups
Readers not familiar with Lie groups can easily verify that set $\{I+UV^{T}|U\in\mathbb{R}^{n\times r},\,\det(I+V^{T}U)\neq 0,V{\rm\,is\,fixed}\}$ along with the matrix multiplication operation forms a Lie group by showing that it includes the identity matrix (when $U=0$), each of its elements has an inverse when $\det(I+V^{T}U)\neq 0$, and lastly matrix multiplication is associative. Similarly, with a given $U$, set $\{I+UV^{T}|V\in\mathbb{R}^{n\times r},\,\det(I+V^{T}U)\neq 0,U{\rm\,is\,fixed}\}$ forms another Lie group. But, generally set $\{I+UV^{T}|U,V\in\mathbb{R}^{n\times r}\,{\rm and}\,\det(I+V^{T}U)\neq 0\}$ does not form a group since the products of its two elements may no long belong to the same set. Hence, I choose to update either $U$ or $V$, not both, at each step in my implementations. Please check the supplementary materials of [19] for further details.